### v1

# size of isogeny classes of abelian varieties in the extended lubin-tate case

Abstract.
We prove a lower bound for the size of the isogeny class of a simple abelian variety over a finite field with commutative endomorphism ring in the extended Lubin-Tate case. Moreover, based on the expected size of the isogeny classes in the Newton stratum, we conjecture that this lower bound is sharp.
1. Introduction
1.1. The Newton stratum
Throughout this paper, ${\mathbb{F}}_{q}$ will denote the finite field with $q$ elements where $q$ is a power of an odd prime. We will denote the algebraically closed field $\overline{{\mathbb{F}}}_{q}$ by $k$. Let $\mathcal{A}_{g,1}$ denote the moduli space of $g$ dimensional principally polarised abelian varieties in characteristic $p.$ This moduli space is often studied by describing it as a disjoint union of locally closed subvarieties. One way to do this is to introduce a “stratification” using the Newton polygon which is an invariant in characteristic $p$. The Newton stratum $N$ associated to an abelian variety $A,$ consists of the points of $\mathcal{A}_{g,1}$ that have the same Newton polygon as that of $A$. In particular, it contains all the elements in the isogeny class of $A.$ In his paper [Oor04], Frans Oort gives an “almost product” structure to the Newton stratum, which we call a “foliation”. The foliation consists of two “transversal” subvarieties known as the “central leaf” and the “isogeny leaf”. The central leaf contains all the abelian varieties over ${\mathbb{F}}_{q^{n}}$ whose $p$-divisible group is isomorphic to that of $A$ over $k$. In particular, it contains all the abelian varieties that are isogenous to $A$ by prime-to-$p$ isogenies. While the “isogeny leaf” consists of the abelian varieties that are isogenous to $A$ by a certain $p$-power isogeny.
Definition 1.1.
(Isogeny class) Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$. We define $I(A,{\mathbb{F}}_{q^{n}})$, the set of principally polarized abelian varieties $y\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})$ isogenous to $A$ over $k$. We call this the isogeny class of $A$ defined over ${\mathbb{F}}_{q^{n}}.$
1.1.1. Estimating the set $I(A,{\mathbb{F}}_{q^{n}})$
Let $c$ and $I$ denote the central leaf and the isogeny leaf of $A$ respectively. We know that the number of Weil $q^{n}$-numbers with a fixed Newton polygon is roughly $O(q^{n\dim c/2})$. This is computed in [ST18]. As one should expect, the size of $I(A,{\mathbb{F}}_{q^{n}})$ should be $O(q^{n(\dim c/2+\dim I)})$ to account for roughly $q^{n\dim N}$ points of $N({\mathbb{F}}_{q^{n}})$. The dimensions of the leaves can be computed explicitly using the Newton polygon of $A$ as proved by Oort in [Oor09].
1.1.2. Previous cases
The question of counting isogeny classes is trivial for supersingular abelian varieties. Indeed, all the abelian varieties in the supersingular Newton stratum are isogenous over $k$. Hence the Newton stratum is the entire isogeny class in this case. The lower bounds of $I(A,{\mathbb{F}}_{q^{n}})$ for the ordinary and the almost ordinary cases is worked out in [ST18] and [OS20] respectively. In another paper [BF23] with Yu Fu, we estimate isogeny classes of certain abelian varieties with real multiplication. We provide upper and lower bounds in this case. The second author of the paper, has work in this direction as well. For example see [Fu23] for the case of a product of a supersingular elliptic curve with an ordinary one.
In this paper, we further generalise the work for the Newton stratum of extended Lubin-Tate type.
Definition 1.2.
A $p$-divisible group ${\mathscr{G}}$ is of extended Lubin-Tate type if its local-local part (or its dual) is trivial or dimension $1$
In the case of $p$-divisible groups of principally polarised abelian varieties, the non-trivial case corresponds to when the $p$-divisible group decomposes as a dimension one and a co-dimension one $p$-divisible group. We note that in this case, the Newton polygon of ${\mathscr{G}}$ has slope $\frac{1}{g}$ and $\frac{g-1}{g}.$
1.1.3. The main theorem
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a geometrically simple, abelian variety with Weil $q$-number $\alpha$. Moreover, let $A[p^{\infty}]$ be its $p$-divisible group which we assume to be of extended Lubin-Tate type. Furthermore, we assume that $K={\mathbb{Q}}(\alpha)$ is the endomorphism algebra of $A\otimes{\mathbb{F}}_{q^{n}}$ such that $K\otimes{\mathbb{Q}}_{p}$ is unramified. Let $N(A,{\mathbb{F}}_{q^{n}})$ denote the size of the set $I(A,{\mathbb{F}}_{q^{n}})$. In this paper, we prove the lower bound for $N(A,{\mathbb{F}}_{q^{n}})$ given in the following main result of this paper.
Theorem 1.3.
For a positive density set of positive integers $n$, we have the following lower bound
$N(A,{\mathbb{F}}_{q^{n}})\geq q^{n(\frac{(g+1)(g-2)}{4}+1+o(1))}.$
Conjecture 1.4.
For a positive density set of positive integers $n$, we have the following estimate
$N(A,{\mathbb{F}}_{q^{n}})=q^{n(\frac{(g+1)(g-2)}{4}+1)+o(1))}.$
The dimension of the Newton stratum is $\frac{g(g-1)}{2}$ and that of the central leaf $c(A)$ is $\frac{(g+1)(g-2)}{2}$. Therefore, the bound in Theorem 1.3 agrees with the expected estimate.
As mentioned in [ST18], one way to prove the upper bound would be to estimate the class groups of all the endomorphisms rings of the isogeny class. We might do this by finding a parametrisation of non maximal orders of $K$ containing the smallest possible endomorphism ring. This seems out of reach. However, methods involving orbital integrals have provided upper bounds in certain cases as well. See for example [AC02].
1.1.4. Comparison of methods from previous cases
The methods used to find bounds in the cases of ordinary and almost ordinary abelian varieties was a consequence of a classification theorem due to Deligne for ordinary abelian varieties and by Oswal-Shankar in the almost ordinary case. In both the classification results, a crucial step is to find “canonical” CM lift(s) of abelian varieties over ${\mathbb{F}}_{q}$ to characteristic zero. Our proof strategy as described below is different from the previous cases in that we do not use lifting methods but count in characteristic zero and then deduce the count mod $p.$
1.2. Overall strategy
1.2.1. Notation
Let $W({\mathbb{F}}_{q})$ denote the Witt vectors over ${\mathbb{F}}_{q}.$ Let $L=W({\mathbb{F}}_{q})[1/p]$ be the unramified extension of ${\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$, while we let $L^{\prime}$ denote $W(k)[1/p]$, the maximal unramified extension of ${\mathbb{Q}}_{p}.$ Let $v$ be the unramified place of $L^{\prime}$ over $p.$
1.2.2. Strategy
As a first step, we note that there are only finitely many abelian varieties over $L^{\prime}$ that reduce to $A$ modulo $v$. Indeed, the number of such abelian varieties over $L^{\prime}$ with endomorphism algebra $K$, will depend on the number of CM-types of $K$. There are only finitely many such. Furthermore, once we fix a CM type on $K$, we get an injective map from the isogeny class over $L^{\prime}$ to the isogeny class in char $p.$ This is proved in section 3, Proposition 3.1. Therefore, we fix an abelian variety $\widetilde{A}$ over $L$ and a CM-type $(K,\Phi)$ on its endomorphism algebra. Our strategy will be to count the number of abelian varieties over $L^{\prime}$ with the endomorphism ring $R$ contained in $K$ characterised by the smallest local conditions. We compute the local conditions in Section 4. We then restrict our count to the set $I(\widetilde{A},L^{\prime}).$ That is, we count the number of principally polarised abelian varieties isogenous to $\widetilde{A}$ with endomorphism ring $R.$ This is done in section 6 and 8. Note that this count is being done in characteristic 00. By the work of Kisin-Lam-Shankar-Srinivasan [KLSS] we know that the number of abelian varieties over ${\mathbb{F}}_{q^{n}}$ having CM lifts lie in finitely many central leaves. See Theorem $1.3$ in [KLSS] for a precise statement. In fact, in this case taking the special fibre of abelian varieties over $W({\mathbb{F}}_{q^{n}})$ will land in exactly one central leaf (see sub-section 5.2). In order to get the final count, we will need to multiply the count in characteristic zero with $q^{\dim I.}$ It is therefore sufficient to prove that $N(\widetilde{A},L^{\prime})\geq q^{\dim c/2}.$
1.2.3. The unpolarised count
In order to make use of the theory of moduli spaces in characteristic $p$, we need to restrict our question to principally polarised abelian varieties. Indeed, in the polarised case, the completion of $\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{p}$ at a point $A_{0}$, denoted by $\mathcal{A}_{g,1}^{\backslash A_{0}}=\operatorname{spf}(R)$ represents the deformation functor of $A_{0}$ as a formal scheme. If we consider the universal deformation $\mathscr{A_{0}}$ over $\operatorname{spf}(R)$, then that can be algebraised to give us an abelian scheme over $\operatorname{Spec}R$ since $\mathscr{A}$ is principally polarised. The (polarised) central leaf of $A_{0}$ is defined to be all $s\in\operatorname{Spec}(R)$ such that the $p$-divisible group $\mathscr{G}_{s}\otimes k$ of the point is isomorphic to that of $A_{0}$ over $k.$ In the unpolarised case, the universal deformation $\mathscr{A}^{\operatorname{np}}$ over $\operatorname{spf}(S)$ of $A_{0}$ need not algebraise. However, its $p$-divisible group over $\operatorname{spf}(S)$ comes from a $p$-divisible group $\mathscr{G}^{\operatorname{np}}$ over $\operatorname{Spec}(S)$. The notion of (unpolarised) central leaf still makes sense. Suppose $x_{0}$ is the closed point of $\operatorname{Spec}(S)$ that represents $A_{0}$. Then the (unpolarised) central leaf of $A$ is the set of all points $s\in S$ such that $\mathscr{G}_{s}^{\operatorname{np}}\otimes k$ is isomorphic to the $p$-divisible group of the special fibre $\mathscr{G}^{\operatorname{np}}_{x_{0}}$ over $k$. This represents a formal object in the local deformation space of $A_{0}$. The dimension of the unpolarised central leaf is also determined by the Newton polygon of $A_{0}$ and can be computed using the methods outlined in [Oor09].
Now, without referring to moduli spaces, we can ask for the count of isogeny classes of abelian varieties over finite fields without any conditions of polarisation. This count in characteristic zero will be the same as computed in subsection 7.0.2 before we take polarisation into account. This is recorded in corollary 7.2. We note that analogous to the case of polarisation, the lower bound for this estimate is in fact of order $\mathcal{O}(q^{\dim c_{\operatorname{np}}/2})$ where $c_{\operatorname{np}}$ is the unpolarised central leaf in the local deformation space of $A.$ Hence, the count over finite fields will be greater than or equal to $q^{n(\dim(c_{\operatorname{np}}/2+I_{\operatorname{np}})+o(1))}$ where $I_{\operatorname{np}}$ is the unpolarised isogeny leaf of the $p$-divisible group of $A_{0}$.
2. Acknowledgements
I am grateful to my advisor, Ananth Shankar for suggesting this project and for many helpful discussions regarding this paper. I also thank Asvin G. for a helpful conversation regarding Lemma 3.2 of the paper. I was partially supported by the NSF grant DMS-2100436.
3. Injectivity of the special fibre map
From this section onwards, we fix a geometrically simple abelian variety $\widetilde{A}$ over $L$ such that $\operatorname{End}^{0}(A)=K.$ Moreover, we assume that $\widetilde{A}$ has good reduction at $p.$ By choosing an embedding of $L$ into ${\mathbb{C}},$ every such abelian variety in characteristic 00 induces a CM type $\Phi$ on $K$. Since $\Phi$ determines the isogeny class of $\widetilde{A}$ over an algebraically closed field of characteristic 00 (see Proposition 1.5.4.1 in [CCO]), we work with the corresponding CM type on $K.$ As $\widetilde{A}$ is simple, the CM type $\Phi$ is primitive. The purpose of this section is to prove the following proposition.
Proposition 3.1.
Let $\widetilde{A}$ and $\widetilde{B}$ be abelian varieties over $L^{\prime}$ with good reduction and the same CM type $(K,\Phi).$ Let $\mathscr{A}$ and $\mathscr{B}$ denote their Néron models respectively over $W(k).$ We assume that the special fibre of both the schemes is $A.$ Then $\widetilde{A}$ and $\widetilde{B}$ are isomorphic.
The following two results are technical lemmas that will help us build a method to prove the proposition. We put together these results in 3.0.1 to give the overall idea of the proof. The subsequent results will then help us prove the proposition.
Lemma 3.2.
Let $\widetilde{A}$ and $\widetilde{B}$ are abelian varieties defined over $L^{\prime}$ such that they have good reduction at a place $v$ of $L^{\prime}$. Suppose that they induce the same CM-type over the algebraically closed field ¯$\overline{\mathbb{Q}}_{p}$. That is, they are isogenous over ¯$\overline{\mathbb{Q}}_{p}$. Then they are isogenous over $L^{\prime}$.
Proof.
Let $\widetilde{A}$ and $\widetilde{B}$ be isogenous over a finite extension $F^{\prime}$ of $L^{\prime}$. Since $\widetilde{A}$ and $\widetilde{B}$ both have good reduction, we see that $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})\hookrightarrow\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$. Let $I_{\overline{v}}$ denote the inertia group of $v$ where $\overline{v}$ is the extension of the place $v$ to the separable closure of $L^{\prime}$. We note that $I_{\overline{v}}$ acts trivially on $\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$ and hence on $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})$ as well. Therefore, $\widetilde{A}$ and $\widetilde{B}$ are in fact isogenous over $L^{\prime}.$
Lemma 3.3.
Suppose that $\widetilde{A}$ is an abelian variety over $L^{\prime}$ with good reduction. Let $A$ denote its special fibre. Then $\operatorname{End}\widetilde{A}=\operatorname{End}A.$
Proof.
Let $\operatorname{End}(\mathscr{A})=R_{1}$ and $\operatorname{End}(A)=R_{2}$. We know that taking the special fibre is a fully faithful functor from the category of abelian varieties over $W(k)$ to abelian varieties over $k.$ Therefore we have a containment of orders $R_{1}\subset R_{2}$ inside $K$. We need to show that they are equal. Since $\widetilde{A}$ is a lift of $A$ to $L^{\prime}$, by Grothendieck- Messing theory, this corresponds to a lift of the filtration $\ker F\operatorname{mod}p\subset{\mathbb{D}}\operatorname{mod}p$ where ${\mathbb{D}}$ is the Dieudonné module of $A[p^{\infty}]$ and $F$ is the Frobenius. We note that all the endomorphisms of $A$ that lift to $\widetilde{A}$ are the ones that leave $\ker F$ invariant. Let $\widetilde{\beta}\in R_{2}.$ Since $R_{1}$ and $R_{2}$ are orders in $K$, their quotient is finite. Thus $n\widetilde{\beta}\in R_{1}$ for some $n.$ This implies that $\ker F$ is invariant under $n\widetilde{\beta}$. Note that $\ker F$ is invariant under the action of multiplication-by-$n$ map as it acts by multiplying its elements by $n$. This proves that $\widetilde{\beta}\in R_{1}.$
3.0.1. Outline of the proof of Proposition 3.1
Since $\widetilde{A}$ and $\widetilde{B}$ have the same CM type, we know that they are isogenous over ${\mathbb{C}}.$ In fact, by the Lemma 3.2, we have an isogeny $\widetilde{\alpha}:\widetilde{A}\rightarrow\widetilde{B}$ over $L^{\prime}.$ The Néron mapping theorem extends this to a unique $W(k)$-morphism of schemes $\varphi:\mathscr{A}\rightarrow\mathscr{B}$. The map $\varphi$ induces an endomorphism $\alpha$ of the special fibre of $A$, that is $\alpha:A\rightarrow A.$ Moreover, by Lemma 3.3, this lifts to an endomorphism $\psi$ of $\mathscr{A},$ generically which gives us an endomorphism $\widetilde{\beta}:\widetilde{A}\rightarrow\widetilde{A}$. Let $G_{\varphi}$ and $G_{\psi}$ be the kernels of $\phi$ and $\psi$ respectively. These are finite flat group schemes over $W(k)$ that reduce to the kernel of $\alpha$ mod $v$. We embed them into an $N$-torsion group scheme $\mathscr{A}[n]$ which is the kernel of the multiplication by $n$ map on $\mathscr{A}$ for some positive integer $n.$ We will prove that the kernels of the two maps must agree as group schemes. Moreover, since we have the same group scheme that occurs as the kernel of an isogeny between $\widetilde{A}$ and $\widetilde{B}$ and an endomorphism of $\widetilde{A}$, we deduce that $\widetilde{A}$ is in fact isomorphic to $\widetilde{B}.$ To that end we have the following Lemma.
Lemma 3.4.
Let $G_{\varphi}$, $G_{\psi}$ and $\mathscr{A}[n]$ be as described above. Let $H$ denote the special fibre of $G_{\varphi}$, $G_{\psi}\subset A[n]$. Then the group subgroup-scheme $G_{\varphi}$ equals $G_{\psi}$.
In order to prove Lemma 3.4, we will use Fontaine’s results on the classification of finite flat group schemes of $p$-power order over an unramified base. For reference, we state the results that we use below from [Con99]. We note that these result only holds for odd primes and we therefore restrict to that case as well for the results in this paper.
3.1. Classification results for finite flat group schemes of prime power order over an unramified base.
Definition 3.5 (Finite Honda systems).
The category of finite Honda systems over $k$ consists of tuples $(\mathscr{L},\mathscr{M})$ where $\mathscr{L}$ and $\mathscr{M}$ are $W(k)[F,V]$-modules of finite $W(k)-$ length such that the following is satisfied.
  1. (1)
    $\mathscr{L}\subset\mathscr{M}$ is a $W(k)$-submodule of $\mathscr{M}$.
  2. (2)
    The map $\mathscr{L}/p\mathscr{L}\rightarrow F\mathscr{M}/p\mathscr{M}$ is an isomorphism.
  3. (3)
    The Vershiebung $V$ is injective on $\mathscr{L}$.
Proposition 3.6.
(Fontaine, see [Con99] Lemma 1.3) The category of finite Honda systems over $k$ is abelian. For a morphism $\varphi:(\mathscr{L},\mathscr{M})\rightarrow(\mathscr{L},\mathscr{M})$, the kernel $\ker(\varphi)=(\mathscr{L}^{\prime},\mathscr{M}^{\prime})$ where $\mathscr{M}^{\prime}=\ker(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime}=\mathscr{L}\cap\mathscr{M}^{\prime}$ and $\operatorname{coker}(\varphi)=(\mathscr{L}^{\prime\prime},\mathscr{M}^{\prime\prime})$ where ′′$\mathscr{M}^{\prime\prime}=\operatorname{coker}(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime\prime}$ is the image of the composition of maps of $W(k)$-modules $\mathscr{L}\rightarrow\mathscr{M}\rightarrow\mathscr{M}^{\prime\prime}.$
Theorem 3.7.
(Fontaine, see Theorem 1.4 [Con99]) For $p>2$, there is an anti-equivalence between the category of finite flat commutative group schemes $G$ over $W(k)$ of $p$-power order and the category of finite Honda systems over $k$. Under this equivalence, $G$ maps to $(\mathscr{L},\mathscr{M})$ where $\mathscr{M}$ is Dieudonné module of the special fibre $G_{k}$ and $\mathscr{L}$ is a $W(k)$ sub-module of $\mathscr{M}$ satisfying the three properties in Definition 3.5. See, for example [Con99], page $5-6$ for a complete description.
The proof of Lemma 3.4 will now follow by proving the result in the category of finite Honda systems as follows.
Proof of Lemma 3.4.
Since all the group schemes are embedded inside the kernel of the multiplication map $[n]$, it is enough to consider the cases of group schemes with prime to $p$ order and $p$-power order separately. For $(n,p)=1$, the reduction map of an abelian variety over $L^{\prime}$ to its special fibre induces an isomorphism on the $n$-torsion. Therefore, the prime-to-$p$ subgroups lift uniquely. Moving on to the case where $n=p^{r}$, under the anti-equivalence in Proposition 3.7, we associate the tuples $(\mathscr{L},\mathscr{M}),(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ to the group schemes $\mathscr{A}[p^{r}],G_{\varphi}$ and $G_{\psi}$ respectively. Since both the kernels reduce to the same group scheme over $k,$ we see that $\mathscr{M}_{1}=\mathscr{M}_{2}$ as they correspond to the Dieudonné module of the special fibre $T$. Furthermore, the tuples $(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ are co-kernels of $(\mathscr{L},\mathscr{M})$ in the category of finite Honda systems. Therefore, Proposition 3.6 proves that $\mathscr{L}_{1}=\mathscr{L}_{2}.$ This shows that the group schemes $G_{\varphi}$ and $G_{\psi}$ are the same as subgroup-schemes of $\mathscr{A}[p^{r}]$ and hence, the required kernels are the same. ∎
Finally, the proposition follows from the following argument that we give below.
Proof of Proposition 3.1.
Lemma 3.4 shows that $\ker\widetilde{\alpha}=\ker\widetilde{\beta}$. This gives us maps $\widetilde{\varphi}:\widetilde{A}\rightarrow\widetilde{B}$ and $\widetilde{\psi}:\widetilde{A}\rightarrow\widetilde{B}$ such that $\widetilde{\varphi}\circ\widetilde{\alpha}=\widetilde{\beta}$ and $\widetilde{\psi}\circ\widetilde{\beta}=\widetilde{\alpha}$. Then we can check that $\widetilde{\varphi}\circ\widetilde{\psi}$ and $\widetilde{\psi}\circ\widetilde{\varphi}$ act as identity when we compose them with $\widetilde{\beta}$ and $\widetilde{\alpha}$ respectively, thus completing the proof of proposition.
4. Local conditions for the smallest endomorphism ring
We next find the local conditions of the orders that can occur as an endomorphism ring in the isogeny class of $A$ over ${\mathbb{F}}_{q}.$ We recall that the Newton polygon of $A[p^{\infty}]$ has slope $1/g$ and $(g-1)/g$. Moreover, we assume that the prime $p$ splits in ${\mathcal{O}}_{K}$ as ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}.$ Therefore, the endomorphism algebra $K$ of $A$ satisfies the condition that $K\otimes{\mathbb{Q}}_{p}=K_{1}\oplus K_{2}$ where $K_{1}$ and $K_{2}$ are unramified extensions of ${\mathbb{Q}}_{p}$ of degree $g$.
4.1. Local conditions at the prime $p$
Proposition 4.1.
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomorphism algebra $K$ as described above. Let $\operatorname{End}(A)=R.$ Then the order $R$ is maximal at $p$. More precisely, $R\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$ where ${\mathcal{O}}_{K_{1}}$ (respectively ${\mathcal{O}}_{K_{2}}$) are the maximal orders in $K_{1}$ (respectively $K_{2}$).
Proof.
Let ${\mathbb{D}}$ denote the Dieudonné module of $A[p^{\infty}]$. Note that ${\mathbb{D}}$ admits a decomposition locally as a direct sum ${\mathbb{D}}\otimes{\mathbb{Q}}_{p}={\mathbb{D}}_{1}\times{\mathbb{D}}_{2}$ where ${\mathbb{D}}_{1}$ and ${\mathbb{D}}_{2}$ are modules over $K_{{\mathfrak{p}}_{1}}$ and $K_{{\mathfrak{p}}_{2}}$ respectively. This further gives us a decomposition of the $p$-divisible group $A[p^{\infty}]\simeq{\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ over ${\mathbb{F}}_{q}$ where ${\mathscr{G}}_{1}$ and ${\mathscr{G}}_{2}$ have slopes $1/g$ and $(g-1)/g$ respectively (see Theorem 9.2 in [Oor07]). Since ${\mathscr{G}}_{1}$ cannot be isogenous to ${\mathscr{G}}_{2}$, we see that $\operatorname{End}({\mathscr{G}}_{1}\times{\mathscr{G}}_{2})\simeq\operatorname{End}({\mathscr{G}}_{1})\times\operatorname{End}({\mathscr{G}}_{2})$. We know that Tate’s theorem states that $R\otimes{\mathbb{Z}}_{p}\simeq\operatorname{End}A[p^{\infty}]$. Therefore, it is sufficient to prove that $\operatorname{End}({\mathscr{G}}_{1})$ and $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. This can be shown analogously to an argument in [OS20], Proposition $2.1$ which we give below. Let ${\mathscr{G}}^{\prime}$ be a $p$-divisible group isogenous to ${\mathscr{G}}_{1}$ such that $\operatorname{End}({\mathscr{G}}^{\prime})$ is maximal in its field of fractions. Let $j:{\mathscr{G}}_{0}\rightarrow{\mathscr{G}}_{1}$ be any arbitrary isogeny. Since ${\mathscr{G}}_{1}$ is isomorphic to a quotient of ${\mathscr{G}}_{0}$ by the kernel of $j$, we will show that $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves $\ker j,$ proving that $\operatorname{End}({\mathscr{G}}^{\prime})\subset\operatorname{End}({\mathscr{G}}_{1}).$ Since ${\mathscr{G}}_{0}$ is a one dimensional $p$-divisible group, it has a unique subgroup of order $p^{r}$ for each $r\geq 1$. Therefore, $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves the kernel of $j$. The same argument with the dual of ${\mathscr{G}}_{1}$ proves that $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. ∎
4.2. Local conditions away from $p$
The above proposition proves that, if $R$ is an order that occurs as an endomorphism ring of an abelian variety isogenous to $A$, then $R$, must be maximal at $p.$ Consequently, it differs from the maximal order or more generally, any order that is maximal at $p$ at only finitely many primes away from $p.$ We know that since $R$ contains the Frobenius endomorphism, hence the local order ${\mathbb{Z}}_{l}[\alpha]\subset R_{l}$. Therefore, in order for $R$ to have the “smallest” local conditions, we will require ${\mathbb{Z}}_{l}[\alpha]=R_{l}.$ The following Lemma proves that we can always adjust the local conditions of an order at $l\neq p$ up to isogeny. Thus, it guarantees the existence of an abelian variety in the isogeny class over ${\mathbb{F}}_{q}$ with endomorphism ring $R$ with the local conditions specified at primes other than $p$.
Lemma 4.2.
Let $R$ be an order in $K$ containing ${\mathbb{Z}}[\alpha].$ We fix an endomorphism algebra $K$ for an isogeny class of abelian varieties. Let $A$ over ${\mathbb{F}}_{q}$ be an element of the isogeny class such that $\operatorname{End}(A)=R^{\prime}$. Then we can find an abelian variety $A^{\prime}$ isogenous to $A$ over ${\mathbb{F}}_{q}$ such that $\operatorname{End}(A^{\prime})\otimes{\mathbb{Z}}_{l}=R_{l}$ for all $l\neq p.$
The proof is already given in [Wat69]. But for the sake completeness we present it here as well. We denote by $T_{l}(A)$ the Tate module of $A$ and let $V_{l}(A)=T_{l}(A)\otimes{\mathbb{Q}}_{l}.$
Proof.
Without loss of any generality we assume $R^{\prime}\subset R$. For primes larger than the index of $R^{\prime}$ in $R$, the two orders are locally the same. Let $l_{1},\dots,l_{m}$ be finitely many primes where the two orders are not the same. According to Tate’s theorem $R^{\prime}\otimes{\mathbb{Z}}_{l}$ takes the lattice $T_{l}(A)$ to $T_{l}(A)$ in $V_{l}(A)$ for all primes $l.$ Since $K$ is commutative, $V_{l}(A)$ is a rank one module over $K\otimes{\mathbb{Q}}_{l}.$ Therefore, for any order in $K$, there exists a lattice with that prescribed order as its endomorphism ring. Let $T_{l_{1}}$ be a sub-lattice of $T_{l}(A)$ that is preserved by the order $R\otimes{\mathbb{Z}}_{l_{1}}$. The quotient of the two lattices corresponds to a finite index subgroup of $A.$ We quotient out $A$ by it to get an abelian variety $A^{\prime}$ isogenous to $A$ such that $T_{l_{1}}(A^{\prime})=T_{l_{1}}(A).$ Now repeat this process for all the other primes to finish the proof. ∎
Remark 4.1.
We would like to clarify that from the above Lemma, it would seem that $\operatorname{End}A$ only depends on its localisation at the primes. However, as Waterhouse remarks in [Wat69], the isomorphism type of the order is not determined by the local conditions. That is, we could have two or more orders that are isomorphic locally but may not be isomorphic as orders in $K.$
We summarise the local conditions in the following notation that characterises the orders in endomorphism ring of the isogeny class with “smallest possible local conditions”.
Notation 4.3.
We define $R_{n}\subset K$ to be the order satisfying the following conditions.
  1. (1)
    The order $R_{n}$ contains the order ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$
  2. (2)
    It is maximal at $p,$ that is, $R_{n}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$
  3. (3)
    At the prime $l,$ we have $R_{n}\otimes{\mathbb{Z}}_{l}={\mathbb{Z}}_{l}[\alpha].$
5. Properties of abelian varieties with endomorphism ring $R_{n}$.
In this section, we note certain useful properties of abelian varieties with endomorphism by $R_{n}$. The properties that we prove below are entirely determined by the endomorphism ring $R_{n}$ and will be useful for us in subsequent sections.
5.1. Field of definition
Lemma 5.1.
(First property) Let $B$ be an abelian variety over $k$. Let $R_{n}$ be its the endomorphism ring. Then the abelian variety $B$ is defined over ${\mathbb{F}}_{q^{n}}$.
Proof.
Without any loss of generality, we prove this result for $n=1.$ We write $R_{1}=R.$ We can assume that $B$ is isogenous to an abelian variety $B^{\prime}$ that has endomoprhism by the maximal order so that $R\subset\operatorname{End}B^{\prime}$. Let $\varphi:B^{\prime}\rightarrow B$ be such an isogeny. Therefore the kernel of $\varphi$ is invariant under $R$, in particular it is invariant under the action of $\alpha$, the $q$-Frobenius of $B$. This implies that kernel is also Galois invariant under $\operatorname{Gal}(k/{\mathbb{F}}_{q})$ because it acts by the Frobenius of $B$. This implies that $B$ is in fact defined over ${\mathbb{F}}_{q}.$
5.2. Central leaf
In this sub-section, we will see that the property of $R_{n}$ being maximal at $p$ determines the isomorphism type of the $p$-divisible group of the abelian variety over $k.$ In order to show this, we first state some definitions and results we need.
Notation 5.2.
Let $R_{k}$ denote the ring $W(k)[F,V]$. Under the contravariant Dieudonné theory, we write ${\mathscr{G}}_{n,m}$ to be the $p$-divisible group over $k$ corresponding to the Dieudonné module $R_{k}/R_{k}(F^{n}-V^{m})$, which we denote by ${\mathbb{D}}(G_{m,n})$
We note that under this characterisation, the $p$-divisible group ${\mathscr{G}}_{m,n}$ will have dimension $n$ and height $m+n.$ The following result is well known, but for the sake of completeness, we give a proof here.
Lemma 5.3.
There exists a unique $p$-divisible group of dimension one (and co-dimension one) over $k.$
Proof.
Let ${\mathscr{G}}$ be a $p$-divisible group of dimension one and height $n$ over $k$. let ${\mathbb{D}}$ be its Dieudonné module. We prove the proposition for dimension one case. The case of co-dimension will follow from the uniqueness of ${\mathscr{G}}$ and the fact that ${\mathscr{G}}_{1,n}^{\vee}$, the dual ${\mathscr{G}}_{1,n}$ is ${\mathscr{G}}_{n,1}$. We know that ${\mathscr{G}}$ is isogenous to ${\mathscr{G}}_{1,n}$ over $k$. We need to show that in fact, they are isomorphic. We first note that ${\mathbb{D}}$ and ${\mathbb{D}}({\mathscr{G}}_{1,n})$ have the same rank, that is they are both of rank $n+1$ as $W(k)$-modules. We only need to show that they are isomorphic as $R_{k}$-modules. We show this by constructing a suitable $R_{k}$ bases for ${\mathbb{D}}$. Note that ${\mathbb{D}}/F{\mathbb{D}}$ is a one dimensional vector space over $k$, and therefore, spanned by (the image of) an element, say $e_{1}$ in ${\mathbb{D}}.$ Now, let $Fe_{1}=e_{2}.$ We note that $Ve_{2}=pe_{1}.$ Next, we let $Fe_{2}=e_{3}$. Note that $e_{3}$ cannot be in the span of $e_{2}$, otherwise $e_{2}$ would be in the span of $e_{1}.$ Moreover, $e_{3}$ cannot be the $W(k)$-span of $e_{1}$ and $e_{2}.$ For if $e_{3}=Fe_{2}=ae_{1}+be_{2}=ae_{1}+bFe_{1}$, then $ae_{1}=0$ in ${\mathbb{D}}/F{\mathbb{D}}$. This implies that $a=0$ which is not possible. We carry on the same process until we hit a bases such that $F$ acts as $Fe_{i}=e_{i+1}$ for $1\leq i\leq n$. Finally writing $Fe_{n+1}=a_{1}e_{1}+a_{2}e_{2}+\dots+a_{n+1}e_{n+1}$ and applying $V$ on both sides implies that $pe_{n+1}=a_{1}Ve_{1}$ and $a_{i}=0$ for $i\geq 2$. Since $e_{1}$ is not in the image of $F,$ we see that $a_{1}=p$ and hence $Fe_{n+1}=pe_{1}$. This gives us an isomorphism between ${\mathscr{G}}$ and ${\mathscr{G}}_{1,n}$.
Definition 5.4.
(Central leaves) Let ${\mathscr{G}}$ be a $p$-divisible group over ${\mathbb{F}}_{q}.$ Then the central leaf of ${\mathscr{G}}$ in $\mathcal{A}_{g,1}\otimes{\mathbb{F}}_{q}$ is defined as
$c({\mathscr{G}})=\{B\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})\mid{\mathscr{G}}\otimes k\simeq B[p^{\infty}]\otimes k\}.$
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a principally polarised abelian variety. Then the central leaf $c(A)$ is defined to be the central leaf associated to its $p$-divisible group $A[p^{\infty}]$.
Lemma 5.5.
(Second property) Let $B$ be an abelian variety over ${\mathbb{F}}_{q^{n}}$ with endomorphism ring $R_{n}\subset K$. Let $c$ be the central leaf of ${\mathscr{G}}_{1,g}\times{\mathscr{G}}_{g-1,g}$. Then $B$ lies in the central leaf $c.$
Proof.
We know that $\operatorname{End}(B[p^{\infty}])\otimes k={\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$. This implies that its Dieudonné module ${\mathbb{D}}\otimes k$ splits into a direct sum ${\mathbb{D}}_{1}\oplus{\mathbb{D}}_{2}$ such that $\operatorname{End}{\mathbb{D}}_{1}$ and $\operatorname{End}{\mathbb{D}}_{2}$ are maximal at $p.$ This gives the desired decomposition of $B[p^{\infty}]\otimes k$ into a product ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$. We know that ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ is isogenous to ${\mathscr{G}}_{1,g}\times{\mathscr{G}}_{g-1,g}$. But since the endomorphism ring preserves the kernel, it must also factor into a direct summand. We see that ${\mathscr{G}}_{1}$ (respectively ${\mathscr{G}}_{2}$) is isogenous, hence isomorphic to ${\mathscr{G}}_{1,g}$ (respectively ${\mathscr{G}}_{g-1,g}$). This shows that $B$ lies in the central leaf $c.$
6. Isogeny classes and the class group
Let $R_{n}\subset K$ be the order defined in Notation 4.3. Let $I\subset K$ be an invertible $R_{n}$ module, that is, $I$ is a fractional ideals of $R_{n}$ such that there exists some fractional ideal $J$ with $IJ=R_{n}$. We say two fractional ideals, $I_{1}$ and $I_{2}$ are equivalent if and only if there exists some $a\in K^{\times}$ such that $I_{1}=aI_{2}.$
Definition 6.1.
(Class group of an order) Under the equivalence mentioned above, we define the equivalence classes of fractional ideals of $R_{n}$ to be the class group of $R_{n}$, denoted as $cl(R_{n}).$
The heart of rest of the paper lies in the following proposition that gives us a convenient characterisation of (isogenous) abelian varieties with endomorphism ring $R_{n}$ over $L^{\prime}.$
Proposition 6.2.
Let $\Phi$ be the primitive CM type on $K$ corresponding to $\widetilde{A}.$ Then there exists an injective map from the class group of $R_{n}$ to the set of isomorphism classes of abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ inside $(K,\Phi).$
Proof.
Let $I$ be an element of $\operatorname{cl}(R_{n}).$ We recall that $\Phi$ gives is an isomorphism $K\otimes{\mathbb{C}}={\mathbb{C}}^{\Phi}$ where ${\mathbb{C}}^{\Phi}$ denotes $g$ copies of ${\mathbb{C}}$ indexed by the embeddings in $\Phi.$ Since $I$ is a lattice in $K,$ we can embed it inside ${\mathbb{C}}^{\Phi}$. We refer to this embedding as $\Phi(I).$ The complex torus ${\mathbb{C}}^{\Phi}/\Phi(I)$ admits a polarisation (see sub-section 8) and therefore we get an abelian variety over ${\mathbb{C}}$ which we denote by $A_{\Phi(I)}$. Since $\Phi$ is primitive, $A_{\Phi(I)}$ is simple and so its endomorphism algebra is $K$. By construction $\operatorname{End}(A_{\Phi(I)})=R_{n}.$ Furthermore, since $A_{\Phi(I)}$ is a CM abelian variety, it descends, with its entire endomorphism algebra to a number field. Now, if the abelian variety had CM by the maximal order, then the abelian variety is defined over the maximal unramified extension of the reflex field $K^{*}$ and hence it would be defined over an unramified extension of ${\mathbb{Q}}_{p}.$ In this case, however, the field of definition might be slightly larger than that and might possibly be a ramified extension of ${\mathbb{Q}}_{p}$. We first rule out this possibility for the abelian variety $A_{\Phi(R_{n})}.$ We note that $A_{\Phi(R_{n})}$ is isogenous to the abelian variety $A^{\prime}$ corresponding to the lattice ${\mathcal{O}}_{K}$ in $K$. Now, $A^{\prime}$ has CM by a maximal order over ¯$\overline{{\mathbb{Q}}}_{p}.$ Let $F$ be the unramified extension of ${\mathbb{Q}}_{p}$ where $A^{\prime}$ is defined. Since the kernel of the isogeny has order equal to the index of $R_{n}$ in ${\mathcal{O}}_{K}$, we can embed it inside $A^{\prime}[m]$ for some $(m,p)=1$. Now the kernel is defined over some extension $F^{\prime}$ of $F$. This must be unramified because the inertia group $I(\overline{v})$ acts trivially on the $m$-torsion and therefore on the subgroup as well. Here $\bar{v}$ refers to the extension of a place $v$ of $F^{\prime}$ to its separable closure. This descends $A_{\Phi(R_{n})}$ to an unramified extension of ${\mathbb{Q}}_{p}.$ This same argument works in descending an abelian variety corresponding to a fractional ideal $I$ whose index in $R_{n}$ is prime to $p.$ So suppose that $P$ is a fractional ideal of $R_{n}$ with index, a power of $p$. We extend this to a fractional ideal $P^{\prime}$ of ${\mathcal{O}}_{K}$. Since $R_{n}$ is maximal at $p$, the index of $P$ in $R_{n}$ equals the index of $P^{\prime}$ in ${\mathcal{O}}_{K}$. Hence the index of $P$ in $P^{\prime}$ is prime to $p.$ Since the abelian variety corresponding to $P^{\prime}$ has endomorphism by the maximal order, it descends to an unramified extension of ${\mathbb{Q}}_{p}.$ We can now repeat above argument with $A_{\Phi(P)}$. Finally, under the equivalence of ideals in $cl(R_{n})$, two equivalent ideals would correspond to isomorphic lattices and hence give us isomorphic abelian varieties. ∎
We note that all the abelian varieties constructed in 6.2 are isogenous to $\widetilde{A}_{L^{\prime}}$ over $L^{\prime}$ Lemma 3.2. Moreover, their special fibres are in fact defined over ${\mathbb{F}}_{q^{n}}$ by Lemma 5.1 and have $\alpha^{n}\in K$ as their Frobenius morphism.
7. Estimates
The goal of this section is to find the asymptotics for $cl(R_{n})$. By proposition 6.2, this gives a lower bound for $N(A,q^{n}).$ We prove the following estimate.
Proposition 7.1.
For a density one set of positive integers, the lower bound for $cl(R_{n})$ is given by $q^{n{(g^{2}-2g/2)}}.$
Proof.
The class group $\operatorname{cl}(R_{n})$ is approximated the square root of the discriminant of $R_{n}$. We first write down the discriminant of ${\mathbb{Z}}[\alpha].$ Let $\alpha^{\prime}_{i}$, where $1\leq i\leq 2g$ be the conjugates of $\alpha.$ For each $i$, write $\alpha^{\prime}_{i}=q^{1/2}(\cos{\theta}+i\sin{\theta}).$ Then discriminant of ${\mathbb{Z}}[\alpha^{n}]$ is given by
$\prod_{i<j\leq 2g}({\alpha^{\prime}_{i}}^{n}-{\alpha^{\prime}_{j}}^{n})^{2}=q^{\frac{2ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{i})-\cos(n\theta_{j})-i\sin(n\theta_{j}))^{2}$
Next, we estimate the index of ${\mathbb{Z}}[\alpha^{n}]$ in $R_{n}.$ The $\operatorname{disc}R_{n}$ is then estimated by dividing the two estimates. Note that the global index is the product of the local indices. Since the index is $1$ at all primes $l\neq p$, it is sufficient to find a bound for the index after tensoring the orders with ${\mathbb{Z}}_{p}$. We further split the calculation in two parts as follows.
7.0.1. Local analysis.
We make the argument for $n=1$ and the general result follows similarly. Under the embedding of ${\mathbb{Z}}[\alpha]\otimes{\mathbb{Z}}_{p}$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}},$ the image of $\alpha$ is $(\alpha_{1},\alpha_{2})$ where $v_{{\mathfrak{q}}}(\alpha_{1})=\frac{1}{g}$ and $v_{q}(\alpha_{2})=\frac{g-1}{g}$. Locally we are reduced to finding the index of ${\mathbb{Z}}_{p}[(\alpha_{1},\alpha_{2})]$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$. The ring ${\mathbb{Z}}_{p}[(\alpha_{1},\alpha_{2})]$ is monogenic and is isomorphic to ${\mathbb{Z}}_{p}[x]/(f_{1}f_{2})$ where $f=f_{1}f_{2}$ is the factorization of the minimal polynomial $f$ of $\alpha$ in ${\mathbb{Z}}_{p}.$ To find the discriminant of the element, we need to take the difference of the roots of $f_{1}f_{2}$. Let $\beta_{i}$ and $\gamma_{i}$ be the roots of $f_{1}$ and $f_{2}$, that is, the conjugates of $\alpha$. We first divide (the image of) $\alpha$ in ${\mathcal{O}}_{K_{1}}\oplus{\mathcal{O}}_{K_{2}}$ by an appropriate power of $q$ so that the valuation becomes 00. The new element, which we call by $\beta$, gives a monogenic ring ${\mathbb{Z}}_{p}[\beta]$ that is isomorphic to ${\mathbb{Z}}_{p}[x]/g_{1}g_{2}.$ The roots of $g_{1}$ and $g_{2}$ are $\beta_{i}/q^{1/g}$ and $\gamma_{i}/q^{g-1/g}$ for $1\leq i\leq g.$ Thus, we have ${\mathbb{Z}}_{p}[\alpha]\subset{\mathbb{Z}}_{p}[\beta]\subset{\mathcal{O}}_{K}.$ The index $[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]$ is computed using the following relation.
$\operatorname{disc}{\mathbb{Z}}_{p}[\beta]\cdot[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\alpha]$
Up to a unit, the index is given by the following expression.
$\frac{\prod_{1\leq i<j\leq g}(\beta_{i}-\beta_{j})(\gamma_{i}-\gamma_{j})(\beta_{i}-\gamma_{j})}{\prod_{1\leq i<j\leq g}(\beta_{i}/q^{1/g}-\beta_{j}/q^{1/g})(\gamma_{i}/q^{(g-1)/g}-\gamma_{j}/q^{(g-1)/g)})(\beta_{i}/q^{1/g}-\gamma_{j}/q^{g-1/g})}$
$=(q^{1/g})^{g(g-1)/2}\cdot(q^{(g-1)/g})^{g(g-1)/2}\cdot(q^{1/g})^{g^{2}}=q^{\frac{g^{2}+g}{2}}$
We next show that the index ${\mathbb{Z}}_{p}[\beta^{n}]\subset{\mathcal{O}}_{K}$ stays constant for most $n$. We consider ${\mathbb{Z}}_{p}[\beta]\subset{\mathcal{O}}_{K}.$ We need to show that for most $n$, the index of ${\mathbb{Z}}_{p}[\beta]$ is the same as that of ${\mathbb{Z}}_{p}[\beta^{n}].$ We do this by computing the discriminant of $\beta^{n}$ and showing that it’s the same for most $n$. We denote by $\omega_{i}$ the conjugates of $\beta$, that is the roots of $g_{1}$ and $g_{2}$ as described above.
$\operatorname{disc}{\mathbb{Z}}_{p}[\beta^{n}]=\prod_{1\leq i<j\leq 2g}(\omega_{i}^{n}-\omega_{j}^{n})^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\beta]\cdot\prod_{1\leq i<j\leq 2g}(\omega_{i}^{n-1}+\omega_{i}^{n-2}\omega_{j}+\dots+\omega_{j}^{n-1})$
Now, note that the residue field of ${\mathcal{O}}_{K_{1}}$ (and ${\mathcal{O}}_{K_{2}}$) is ${\mathbb{F}}_{p^{g}}.$ So if we take $n$ larger than $p^{g}-1$ then all the conjugates of $\beta$ are $1\operatorname{mod}$ $p$. Moreover, if we take $n$ co-prime to $p$, then the discriminant of ${\mathbb{Z}}_{p}[\beta^{n}]$ and ${\mathbb{Z}}_{p}[\beta]$ differ by a unit. This proves that the index of ${\mathbb{Z}}[\beta^{n}]\subset R_{n}$ stays the same for a density one set of positive integers $n$.
7.0.2. Back to the global setting.
Up to some constant, the discriminant of $R_{n}$ is bounded above by the following expression.
$\frac{q^{\frac{ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{j})-\cos(n\theta_{j})-\sin(n\theta_{j}))^{2}}{q^{n(g^{2}+g)}}$
Since the term with sines and cosines is bounded, we see that the class group $cl(R_{n})$ should be approximated by $q^{\frac{2ng(2g-1)}{2}-\frac{n(g^{2}+g)}{2}}=q^{n(g^{2}-2g)/2}$
Corollary 7.2.
For a positive density set of positive integers $n$, the size of (unpolarised) isogeny class of $\widetilde{A}$ in characteristic zero is greater than or equal to $q^{n(\frac{(g^{2}-2g)}{2}+o(1))}$.
8. Polarisation
So far we have computed a lower bound for the number of isomorphism classes of simple abelian varieties with endomorphism ring $R_{n}$ and CM-type $\Phi.$ In order to restrict this count to the isogeny class in $\mathcal{A}_{g,1},$ we need to take polarisation into account. Howe in [How95] describes the polarisation data for ordinary abelian varieties which is summed up nicely in [ST18], section 3.1. However, since the data is described purely in characteristic zero and is independent of the type of abelian variety, we use the same description in our case.
8.1. Polarisation data.
We give a brief summary of the description of polarisation on . We fix a CM type $\Phi$ on $K$. The polarisation on an abelian variety corresponding to $I$ in $cl(R_{n})$ is given by a $K$-bilinear form on $K$
$\Psi:(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(\lambda x\overline{y})$
with the following properties to keep in mind.
  1. (1)
    Here, $\lambda\in K$ is purely imaginary.
  2. (2)
    For all $\varphi\in\Phi$, we have $\varphi(\lambda)/i$ is positive.
  3. (3)
    The dual of an ideal $I$ is the trace dual of $\overline{I}$ where $\overline{I}$ is the CM conjugate of $I$ with respect to the unique involution on $K.$
We note that if an ideal $I$ is dual to itself with respect to the above form then then its corresponding abelian variety will be principally polarized. To restrict our count to principally polarised abelian varieties, we have the following proposition. The proof of this proposition goes through analogously to Lemma 3.8 in [ST18].
Proposition 8.1.
Let $R_{n}^{+}$ denote the intersection of $R_{n}$ with the maximal totally real field $K^{+}\subset K.$ Then the kernel of the norm map
$N:cl(R_{n})\rightarrow cl^{+}(R_{n}^{+})$
is either empty or is in one-to-one correspondence with principally polarised abelian varieties isogenous to $A$ over ${\mathbb{F}}_{q^{n}}.$ Here $cl^{+}(R_{n}^{+})$ denotes the narrow class group of the order $R_{n}^{+}$ and the norm map refers to $I\mapsto I\overline{I}$.
Lemma 8.2.
For a positive density set of integers $n,$ we have $\frac{|cl(R_{n})|}{|cl^{+}(R_{n}^{+})|}=q^{\frac{n(g+1)(g-2)}{2}}$
Proof.
It suffices to find an upper bound for $cl(R_{n}^{+})$. Using a similar strategy from the previous computation, we will bound the index $[{\mathbb{Z}}[\alpha_{n}^{+}]:R_{n}^{+}].$ For simplicity we do the computation with $n=1.$ Let $\alpha^{+}$ be the element $\pi+q/\pi$. We denote the $p$-adic conjugates of $\alpha^{+}$ by $\alpha_{i}^{+}$ for $1\leq i\leq g.$ Following the same strategy as the previous section, the index of ${\mathbb{Z}}[\alpha^{+}]$ inside $R_{n}^{+}$ estimated by their index at $p$. The index is given by the following expression at $p$.
$\frac{\prod_{1\leq i<j\leq g}(\alpha_{i}^{+}-\alpha_{j}^{+})}{\prod_{1\leq i<j\leq g}(\alpha_{i}/q^{1/g}-\alpha_{j}/q^{1/g})}=q^{\frac{{g-1}}{2}}$
Dividing the discriminant of ${\mathbb{Z}}[\alpha^{+}]$ calculated as in [ST18] with the index obtained above, we get that $cl(R^{+})$ is approximated by $q^{\frac{g^{2}-3g+2}{4}}$. Combining this with the estimate for $cl(R_{n})$ we get the desired result. ∎
We note that if $\iota:\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ is a polarisation on $\widetilde{A}$, then it extends to a polarisation on its Néron model $\mathscr{A}$ and hence, reduces to a polarisation on the special fibre as well. Therefore, the count in characteristic zero restricted to principally polarised varieties gives us the right estimate in characteristic $p$. It remains to prove that the kernel of the Norm map in Proposition 8.1 is non empty. That is, the subset of $I(\widetilde{A},L^{\prime})$ with endomorphism ring $R_{n}$ is not empty. This is proved in the final proposition of this paper.
Proposition 8.3.
We assume that $q^{1/2},\alpha=\alpha_{1},$ and its conjugates $\alpha_{2},\dots\alpha_{g}$, are multiplicatively independent. Then there exist principally polarised abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ for a positive density set of integers $n.$
Proof.
Let $R^{\prime}_{n}={\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and let $h_{n}$ denote the minimal polynomial of $\alpha^{n}+(q/\alpha)^{n}$ and let $\lambda_{n}=(\alpha^{n}-(q/\alpha)^{n}h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n}))^{-1}.$ By Proposition 9.4 in [How95], the trace dual of $R^{\prime}$ is $\lambda_{n}^{-1}R^{\prime}.$ Therefore, if we consider the bilinear form
$(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(\lambda_{n}x\overline{y})$
then the dual of $R^{\prime}_{n}$ with respect to the form is itself. Furthermore, the above form induces a polarisation on the elements of $cl(R_{n})$ when $\varphi(\lambda_{n})/i$ is a positive real number for every $\varphi\in\Phi.$ The argument given in Proposition $3.6$ of [ST18] shows that for a positive proportion of $n$, $\varphi(\lambda_{n})/i$ is positive under the assumption that the set $\{q^{1/2},\alpha_{1},\dots,\alpha_{g}\}$ is multiplicatively independent. We assume that $n$ is an integer such that $\lambda_{n}$ satisfies appropriate conditions so that we get a polarisation on the elements of $cl(R_{n}).$ Let ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}$ be prime ideals in ${\mathcal{O}}_{K}$ lying over $p.$ Without loss of any generality, we assume $n=1$. Since $R=R_{1}$ is maximal at ${\mathfrak{p}}_{1}$, the ideal $R\cap{\mathfrak{p}}_{1}$ is invertible. The same holds for ${\mathfrak{p}}_{2}$. By abuse of notation we will use the same notation for the $R$-ideals $R\cap{\mathfrak{p}}_{1}$ and $R\cap{\mathfrak{p}}_{2}.$ In the argument that follows, we will go back and forth between abelian varieties and their corresponding lattice in $K$ using Proposition 6.2. We know that the lattice $R_{n}$ of $\widetilde{A}$ over ${\mathbb{Z}}_{p}$ admits a decomposition corresponding to the slope $1/g$ and $g-1/g$ of (the special fibre of) $\widetilde{A}.$ We write this decomposition as $T_{1/g}\oplus T_{g-1/g}.$ We recall that $T_{1/g}$ (and $T_{g-1/g}$) has endomorphism by the maximal order in $K_{1}$ (and $K_{2}$). Now, $R^{\prime}$ and $R$ agree everywhere except at ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2},$ and hence the same is true for $R^{\prime}$ and the dual of $R$. Therefore, the dual of $R$ is ${\mathfrak{p}}_{1}^{m}{\mathfrak{p}}_{2}^{k}$ for some integers $m$ and $k.$ The kernel of the corresponding polarisation map $\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ corresponds to the $\overline{\omega}^{m}$-torsion of $T_{1/g}$ and $\overline{\omega}^{k}$-torsion of $T_{g-1/g}$ where $\overline{\omega}$ is the uniformizer of ${\mathcal{O}}_{K_{1}}$ and ${\mathcal{O}}_{K_{2}}$. Since the polarisation map has a kernel that is self dual, we see that $m=k.$ Therefore, the kernel equals to the $p^{gm}$-torsion of $\widetilde{A}$ and hence $\widetilde{A}$ is principally polarised as $\widetilde{A}^{\vee}\simeq\widetilde{A}/\widetilde{A}[p^{gm}]\simeq\widetilde{A}$.