### v2

# size of isogeny classes of abelian varieties of lubin-tate type

Abstract.
We prove a lower bound for the size of the isogeny class of a simple abelian variety over a finite field with commutative endomorphism ring in the Lubin-Tate case. Moreover, based on the expected size of the isogeny classes in the Newton stratum, we conjecture that this lower bound is sharp.
Key words and phrases:
Abelian varieties, finite fields, complex multiplication, isogeny class, Newton stratum
2000 Mathematics Subject Classification:
11G15 (Primary), 11G10 (Secondary).
1. Introduction
1.1. The Newton stratum
Throughout this paper, we denote ${\mathbb{F}}_{q}$ to be the finite field with $q$ elements where $q$ is a power of an odd prime. For any positive integer $n$, we denote by ${\mathbb{F}}_{q^{n}}$ to be the unique finite field extension of ${\mathbb{F}}_{q}$ of degree $n$. We will denote the algebraically closed field $\overline{{\mathbb{F}}}_{q}$ by $k$. Let $\mathcal{A}_{g,1}$ denote the moduli space of $g$ dimensional principally polarised abelian varieties in characteristic $p.$ This moduli space is often studied by describing it as a disjoint union of locally closed subvarieties. One way to do this is to introduce a “stratification” using the Newton polygon which is an invariant in characteristic $p$. Let $A$ be a principally polarised abelian variety over ${\mathbb{F}}_{q}$ and let $N(A)$ be the associated Newton stratum. We know that this is locally closed subvariety of $\mathcal{A}_{g,1}$. Then $N(A)({\mathbb{F}}_{q^{n}})$ consists of the points of $\mathcal{A}_{g,1}$ defined over ${\mathbb{F}}_{q^{n}}$ that have the same Newton polygon as that of $A$. In particular, it contains all the elements in the isogeny class of $A$ as defined below.
Definition 1.1.
(Isogeny class) We define $I(A,{\mathbb{F}}_{q^{n}})$, to be the set of isomorphism classes of principally polarized abelian varieties $y\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})$ isogenous to $A$ over $k$. We call this the isogeny class of $A$ defined over ${\mathbb{F}}_{q^{n}}.$
In his paper [Oor04], Frans Oort gives an “almost product” structure to the Newton stratum, which we call a “foliation”. The foliation consists of two “transversal” subvarieties known as the “central leaf” and the “isogeny leaf”. The central leaf contains all the abelian varieties over ${\mathbb{F}}_{q^{n}}$ whose $p$-divisible group is isomorphic to that of $A$ over $k$. In particular, it contains all the abelian varieties that are isogenous to $A$ by prime-to-$p$ isogenies. While the “isogeny leaf” consists of the abelian varieties that are isogenous to $A$ by a certain $p$-power isogeny. Using this structure, Shankar and Tsimerman (See [ST18], Section 2) find heuristics for the size of isogeny classes of abelian varieties. We briefly describe this.
1.2. Heuristics for the size of the isogeny class.
Let $c(A)$ and $\iota(A)$ denote the central leaf and the isogeny leaf of $A$ respectively.
Definition 1.2.
(Weil $q$-number of an abelian variety) Let $A$ be as above and $\alpha_{A}$ be its geometric Frobenius. Further assume that $A$ is simple. Then we know that $\alpha_{A}$ is an algebraic integer such that for every embedding $\psi:{\mathbb{Q}}(\alpha_{A})\hookrightarrow{\mathbb{C}}$, we have that $|\psi(\alpha_{A})|=\sqrt{q}.$ Such an algebraic integer is called a Weil $q$-number. By Honda Tate theory we know that the isogeny classes of simple abelian varieties are in bijection with Weil $q$-numbers under $A\mapsto\alpha_{A}$.
We know that the number of Weil $q^{n}$-numbers with a fixed Newton polygon is roughly $O(q^{n\dim c(A)/2})$ which gives us an estimate for the number of ${\mathbb{F}}_{q^{n}}$ isogeny classes in the Newton stratum. This is computed in [ST18], Lemma 2.2. As one should expect, the size of $I(A,{\mathbb{F}}_{q^{n}})$ should be $O(q^{n(\dim c(A)/2+\dim\iota(A))})$ to account for roughly $q^{n\dim N}$ points of $N({\mathbb{F}}_{q^{n}})$. The dimensions of the central leaves and Newton strata can be computed explicitly and only depends on the Newton polygon of $A$. This was also proved by a follow up paper by Oort (See [Oor09], Section 5 and 6).
1.3. Previous work
The question of counting isogeny classes is trivial for supersingular abelian varieties. Indeed, all the abelian varieties in the supersingular Newton stratum are isogenous over $k$. Hence the Newton stratum is the entire isogeny class in this case. One of the first results goes back to the upper bound for elliptic curves by Lenstra [Len87] and subsequent work provides mainly lower bounds in higher dimensions. The lower bounds of $I(A,{\mathbb{F}}_{q^{n}})$ for the ordinary and the almost ordinary cases is worked out in [ST18] and [OS20] respectively. In another paper [BF23] with Yu Fu, we estimate isogeny classes of certain abelian varieties with real multiplication. We provide upper and lower bounds in this case. The second author of the paper, has work in this direction as well. For example, see [Fu23] for the case of a product of a supersingular elliptic curve with an ordinary one.
In this paper, we further generalise the work for the Newton stratum of Lubin-Tate type.
Definition 1.3.
A $p$-divisible group is of Lubin-Tate type if the dimension of its local-local part or that of its dual is atmost one.
In the case of $p$-divisible groups of principally polarised abelian varieties, the non-trivial case corresponds to when the $p$-divisible group decomposes as a dimension one and a co-dimension one $p$-divisible group. We note that in this case, its Newton polygon has slope $1/g$ and $(g-1)/g.$
1.4. The main theorem
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a geometrically simple, abelian variety with Weil $q$-number $\alpha_{A}$. For the sake of convenience, we drop of the subscript and the dependence on $A$ will be implicit. Let the $p$-divisible group $A[p^{\infty}]$ of $A$ be of Lubin-Tate type. Let $K={\mathbb{Q}}(\alpha)$ be the the endomorphism algebra of $A\otimes k$. Due to the condition on the Newton polygon, the tensor product $K\otimes{\mathbb{Q}}_{p}$ is a direct sum of two field extensions of ${\mathbb{Q}}_{p}$ of degree $g$ which we assume to be unramified.
For any positive integer $n$, let $N(A,{\mathbb{F}}_{q^{n}})$ denote the size of the set $I(A,{\mathbb{F}}_{q^{n}})$. We prove the lower bound for $N(A,{\mathbb{F}}_{q^{n}})$ given in the following main result of this paper.
Theorem 1.4.
For a set of positive integers $n$ with positive natural density, we have the following lower bound
$N(A,{\mathbb{F}}_{q^{n}})\geq q^{n(\frac{(g+1)(g-2)}{4}+1+o(1))}.$
Conjecture 1.5.
For a set of positive integers $n$ with positive natural density, we have the following estimate
$N(A,{\mathbb{F}}_{q^{n}})=q^{n(\frac{(g+1)(g-2)}{4}+1+o(1))}.$
The dimension of the Newton stratum is $g(g-1)/2$ and that of the central leaf $c(A)$ is $(g+1)(g-2)/2$. Therefore, the bound in Theorem 1.4 agrees with the expected estimate.
As mentioned in [ST18], one way to prove the upper bound would be to estimate the class groups of all the endomorphisms rings of the isogeny class. We might do this by finding a parametrisation of non maximal orders of $K$ containing the smallest possible endomorphism ring. This seems out of reach. However, methods involving orbital integrals have provided upper bounds in certain cases as well. See for instance [AC02].
1.5. Comparison of methods from previous cases
The methods used to find bounds in the cases of ordinary and almost ordinary abelian varieties was a consequence of a classification theorem due to Deligne for ordinary abelian varieties and by Oswal-Shankar in the almost ordinary case. In both the classification results, a crucial step is to find “canonical” CM lift(s) of abelian varieties over ${\mathbb{F}}_{q}$ to characteristic zero. Our proof strategy as described below is different from the previous cases in that we do not use lifting methods but count in characteristic zero and then deduce the count mod $p.$
1.6. Overall strategy
Let $W(-)$ denote the functor of Witt vectors. Let $L=W({\mathbb{F}}_{q})(1/p)$ be the unramified extension of ${\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$, and $L^{\prime}$ be $W(k)(1/p)$, the maximal unramified extension of ${\mathbb{Q}}_{p}.$ Let $v$ be the unramified place of $L^{\prime}$ over $p.$
Counting in characteristic zero
We count in characteristic zero (over $L^{\prime})$ as follows. We fix $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ as in Theorem 1.4 and an abelian variety $\widetilde{A}$ over $L$ that reduces to $A$ mod $p$. In addition, we fix a CM-type $(K,\Phi)$ on the endomorphism algebra of $\widetilde{A}$.
As a first step, we note that there are only finitely many abelian varieties over $L^{\prime}$ that reduce to $A$ (base changed to $L^{\prime}$) modulo $v$. Indeed, the number of such abelian varieties over $L^{\prime}$ with endomorphism algebra $K$, will depend on the number of CM-types of $K$. There are only finitely many such. Furthermore, once we fix a CM type on $K$, we get an injective map from the isogeny class over $L^{\prime}$ to the isogeny class in char $p.$ This is proved in Section 2, Proposition 2.7. Therefore, it is enough to count the number of abelian varieties over $L^{\prime}$ with a fixed CM type. Let $n$ be any positive integer. Now to get a lower bound on $I(A,{\mathbb{F}}_{q^{n}})$, we reduce the problem to counting the set $I(A,L^{\prime})$ whose endomorphism algebra $K$ contains a certain fixed endomorphism ring $R_{n}$. We then give an injection from $Cl(R_{n})$, the class group of $R_{n}$ to $I(A,L^{\prime})$. We note that the ring $R_{n}$ depends on $n$ and is characterized by certain local conditions at the prime $p$, that we compute in Section 3. It is crucial to take into account these conditions as they help in estimating the class group of $R_{n}$. Moreover, these local conditions ensure that all the abelian varieties in the set $I(A,L^{\prime})$ with endomorphism ring $R_{n}$ descend to $W({\mathbb{F}}_{q^{n}})(1/p)$ (See Lemma 4.1), consequently giving us a sharp lower bound for $I(A,{\mathbb{F}}_{q^{n}})$ once we reduce mod $p.$
Reducing mod $p$
By the work of Kisin-Lam-Shankar-Srinivasan we know that the number of abelian varieties over ${\mathbb{F}}_{q^{n}}$ having CM lifts lie in finitely many central leaves. See Theorem $1.3$ in [KLSS25] for a precise statement. In fact, in this case taking the special fibre of abelian varieties over $W({\mathbb{F}}_{q^{n}})$ will land in exactly one central leaf (See Section 4.2). It is therefore sufficient to prove that $N(\widetilde{A},L^{\prime})\geq q^{\dim c(A)/2}.$ In order to get the final count, we will need to multiply the count in characteristic zero with $q^{\dim\iota(A)}$ thus proving the heuristic in Section 1.2.
1.7. Organisation of the paper
In Section 2 we prove that the reduction mod $p$ map from the isogeny class in characteristic zero to finite fields is injective. In Section 3 and 4, we compute the local conditions for the endomorphism ring that we use to count and certain properties of abelian varieties with that particular endomorphism ring. In Section 6, we give an injection between the isogeny class and the class of the order. In Section 5, we find bounds for the (unpolaised) isogeny class by estimating the class group of the order. Finally, Section 7 is dedicated to taking polarisation into account for the count.
1.8. The unpolarised count
In order to make use of the theory of moduli spaces in characteristic $p$, we need to restrict our question to principally polarised abelian varieties. Indeed, in the polarised case, the completion of $\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{p}$ at an ${\mathbb{F}}_{q}$-point $A$, denoted by $\mathcal{A}_{g,1}^{\backslash A}=\operatorname{spf}(Y)$ represents the deformation functor of $A$ as a formal scheme. If we consider the universal deformation $\mathscr{A}$ over $\operatorname{spf}(Y)$, then that can be algebraised to give us an abelian scheme over $\operatorname{Spec}(Y)$ since $\mathscr{A}$ is principally polarised. The polarised central leaf of $A$ is defined to be all $s\in\operatorname{Spec}(Y)$ such that the $p$-divisible group $\mathscr{G}_{s}\otimes k$ of the point is isomorphic to that of $A$ over $k.$ In the unpolarised case, the universal deformation $\mathscr{A}^{\operatorname{np}}$ over $\operatorname{spf}(X)$ of $A$ need not algebraise. However, its $p$-divisible group over $\operatorname{spf}(X)$ comes from a $p$-divisible group $\mathscr{G}^{\operatorname{np}}$ over $\operatorname{Spec}(X)$. The notion of unpolarised central leaf still makes sense. Suppose $x_{0}$ is the closed point of $\operatorname{Spec}(X)$ that represents $A$. Then the unpolarised central leaf of $A$ is the set of all points $s\in S$ such that $\mathscr{G}_{s}^{\operatorname{np}}\otimes k$ is isomorphic to the $p$-divisible group of the special fibre $\mathscr{G}^{\operatorname{np}}_{x_{0}}$ over $k$. This represents a formal object in the local deformation space of $A$. The dimension of the unpolarised central leaf is also determined by the Newton polygon of $A$ and can be computed using the methods outlined in [Oor09], Section 4. Now, without referring to moduli spaces, we can ask for the count of isogeny classes of abelian varieties over finite fields without any conditions of polarisation. This count in characteristic zero will be the same as computed in subsection 6.2 before we take polarisation into account. This is recorded in corollary 6.2. We note that analogous to the case of polarisation, the lower bound for this estimate is in fact of order $\mathcal{O}(q^{\dim c_{\operatorname{np}}/2})$ where $c_{\operatorname{np}}$ is the unpolarised central leaf in the local deformation space of $A.$ Hence, the count over finite fields will be greater than or equal to $q^{n(\dim(c_{\operatorname{np}}/2+\iota_{\operatorname{np}})+o(1))}$ where $\iota_{\operatorname{np}}$ is the unpolarised isogeny leaf of the $p$-divisible group of $A$.
Acknowledgements
I am grateful to Ananth Shankar for suggesting this project and for many helpful discussions regarding this paper. I also thank Asvin G. for a helpful conversation regarding Lemma 2.1 of the paper. I was partially supported by the NSF grant DMS-2100436.
2. Injectivity of the special fibre map
From this section on wards, we fix $A\in\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{q}$ as in the main Theorem 1.4. Let $\widetilde{A}$ abelian variety $L=W({\mathbb{F}}_{q})(1/p)$ be reducing to $A$ over ${\mathbb{F}}_{q}$. We assume both $A$ and $\widetilde{A}$ to be geometrically simple. Let $\operatorname{End}^{0}(A)=K=\operatorname{End}^{0}(\widetilde{A})$. Choosing an embedding of $L$ into ${\mathbb{C}},$ induces a CM type $\Phi$ on $K$. We fix one such CM type $\Phi$ on $K$ that determines the isogeny class of $\widetilde{A}$ over an algebraically closed field of characteristic 00 (See Proposition 1.5.4.1 in [CCO14]). In fact we have the following lemma:
Lemma 2.1.
Let $\widetilde{A}$ (base changed to $L^{\prime}=W(k)(1/p))$ be as above and $\widetilde{B}$ be another abelian variety defined over $L^{\prime}$ such that they have good reduction at a place $v$ of $L^{\prime}$. Suppose that they induce the same CM-type over the algebraically closed field ¯$\overline{\mathbb{Q}}_{p}$. That is, they are isogenous over ¯$\overline{\mathbb{Q}}_{p}$. Then they are isogenous over $L^{\prime}$.
Proof.
Let $\widetilde{A}$ and $\widetilde{B}$ be isogenous over a finite extension $F^{\prime}$ of $L^{\prime}$. Since $\widetilde{A}$ and $\widetilde{B}$ both have good reduction, we see that $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})\hookrightarrow\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$. Here $\widetilde{A}_{-}$ denotes the abelian variety base changed to the particular field. Let $I_{\overline{v}}$ denote the inertia group of $v$ where $\overline{v}$ is the extension of the place $v$ to the separable closure of $L^{\prime}$. We note that $I_{\overline{v}}$ acts trivially on $\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$ and hence on $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})$ as well. Therefore, $\widetilde{A}$ and $\widetilde{B}$ are isogenous over $L^{\prime}.$
We further observe that the reduction map preserves the endomorphism orders.
Lemma 2.2.
Let $\widetilde{A}$ and $A$ be as above. Then $\operatorname{End}\widetilde{A}=\operatorname{End}A.$
Proof.
Let $\widetilde{\mathscr{A}}$ be the Néron model of $\widetilde{A}$. We denote by $\operatorname{End}(\widetilde{A})=\operatorname{End}(\widetilde{\mathscr{A}})=R$ and $\operatorname{End}(A)=S$. We know that taking the special fibre is a fully faithful functor from the category of abelian varieties over $W(k)$ to abelian varieties over $k.$ Therefore we have a containment of orders $R\subset S$ inside $K$. We need to show that they are equal. Since $\widetilde{A}$ is a lift of $A$ to $L^{\prime}$, by Grothendieck- Messing theory, this corresponds to a lift of the filtration $\ker F\operatorname{mod}p\subset{\mathbb{D}}\operatorname{mod}p$ where ${\mathbb{D}}$ is the Dieudonné module of $A[p^{\infty}]$ and $F$ is the Frobenius. We note that all the endomorphisms of $A$ that lift to $\widetilde{A}$ are the ones that leave $\ker F$ invariant. Let $\widetilde{\beta}\in S.$ Since $R$ and $S$ are orders in $K$, their quotient is finite. Thus $n\widetilde{\beta}\in R$ for some $n.$ This implies that $\ker F$ is invariant under $n\widetilde{\beta}$. Note that $\ker F$ is invariant under the action of multiplication-by-$n$ map as it acts by multiplying its elements by $n$. This proves that $\widetilde{\beta}\in R.$
The main result of this section is Proposition 2.7 which provides an injection between the isogeny class of $\widetilde{A}$ over $L^{\prime}$ to the isogeny class of its reduction mod $p$. The reader may wish to skip to the proposition directly. Its proof involves Fontaine’s results on classification of finite flat group schemes that we first review below.
2.1. Classification of finite flat group schemes of prime power order.
We follow the exposition and results in [Con99] Section 1, and state the results that we apply later. We work over an unramified base and denote by $D_{k}:=W(k)[F,V]$ where $F$ and $V$ will stand for the Frobenius and the Verchiebung. We recall that $p>2$.
Definition 2.3 (Finite Honda systems).
The category of finite Honda systems over $W(k)$ consists of tuples $(\mathscr{L},\mathscr{M})$ where $\mathscr{L}$ and $\mathscr{M}$ are $D_{k}$-modules of finite $W(k)$-length such that the following is satisfied:
  1. (1)
    $\mathscr{L}\subset\mathscr{M}$ is a $W(k)$-submodule of $\mathscr{M}$.
  2. (2)
    The $k$-linear map $\mathscr{L}/p\mathscr{L}\rightarrow\mathscr{M}/F\mathscr{M}$ is an isomorphism.
  3. (3)
    The Vershiebung $V$ is injective on $\mathscr{L}$.
Proposition 2.4.
(Fontaine. See [Con99] Lemma 1.3) The category of finite Honda systems over $k$ is abelian. Let $\varphi:(\mathscr{L},\mathscr{M})\rightarrow(\mathscr{L},\mathscr{M})$ be a morphism in this category. Then the kernel is given by $\ker(\varphi):=(\mathscr{L}^{\prime},\mathscr{M}^{\prime})$ where $\mathscr{M}^{\prime}:=\ker(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime}:=\mathscr{L}\cap\mathscr{M}^{\prime}$. While the cokernel is $\operatorname{coker}(\varphi):=(\mathscr{L}^{\prime\prime},\mathscr{M}^{\prime\prime})$ where ′′$\mathscr{M}^{\prime\prime}:=\operatorname{coker}(\mathscr{M}\rightarrow\mathscr{M})$ and $\mathscr{L}^{\prime\prime}$ is the image of the composition of maps of $W(k)$-modules $\mathscr{L}\rightarrow\mathscr{M}\rightarrow\mathscr{M}^{\prime\prime}.$
Proposition 2.5.
(Fontaine. See Theorem 1.4 [Con99]) For $p>2$, there is an anti-equivalence between the category of finite flat commutative group schemes $G$ over $W(k)$ of $p$-power order and the category of finite Honda systems over $W(k)$.
Under this equivalence, $G$ maps to $(\mathscr{L},\mathscr{M})$ where $\mathscr{M}$ is Dieudonné module of the special fiber which we denote by $G_{k}$ and $\mathscr{L}$ is a $W(k)$ sub-module of $\mathscr{M}$ satisfying the three properties in Definition 2.3. See, for example [Con99], Section $1$ for a summary of Fontaine’s results.
Now, let $\widetilde{A}$ and $\widetilde{B}$ be as in Lemma 2.1. Since $\widetilde{A}$ and $\widetilde{B}$ have the same CM type, by Lemma 2.1, we have an isogeny $\widetilde{\alpha}:\widetilde{A}\rightarrow\widetilde{B}$ over $L^{\prime}.$ The Néron mapping theorem extends this to a unique $W(k)$-morphism between their Néron models $\varphi:\widetilde{\mathscr{A}}\rightarrow\widetilde{\mathscr{B}}$. The map $\varphi$ induces an endomorphism $\alpha$ of the special fibre of $A$, that is $\alpha:A\rightarrow A.$ Moreover, by Lemma 2.2, this lifts to an endomorphism $\psi$ of $\mathscr{A},$ generically which gives us an endomorphism $\widetilde{\beta}:\widetilde{A}\rightarrow\widetilde{A}$. We will need the following lemma.
Lemma 2.6.
We use the notation above. Let $G_{\varphi}$ and $G_{\psi}$ be the kernels of $\varphi$ and $\psi$ respectively. For some $m>0$, we have $G_{\varphi},G_{\psi}\subset\widetilde{\mathscr{A}}[m]$. Let $H\subset A[m]$ denote the special fibre of the two finite flat group schemes. Then $G_{\varphi}$ equals $G_{\psi}$ as subgroup schemes of $\widetilde{\mathscr{A}}[m]$.
Proof.
Let $G_{\varphi}$, $G_{\psi}$ and $\widetilde{\mathscr{A}}[m]$ be as described above. Let $H$ denote the special fibre of $G_{\varphi}$, $G_{\psi}\subset A[m]$. Since all the group schemes are embedded inside the kernel of the multiplication map $[m]$, it is enough to consider the cases of group schemes with prime to $p$ order and $p$-power order separately. For $(m,p)=1$, the reduction map of an abelian variety over $L^{\prime}$ to its special fibre induces an isomorphism on the $m$-torsion. Therefore, the prime-to-$p$ subgroups lift uniquely. For the case where $m=p^{r}$, we associate the tuples $(\mathscr{L},\mathscr{M}),(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ to the group schemes $\widetilde{\mathscr{A}}[p^{r}],G_{\varphi}$ and $G_{\psi}$ respectively, under the anti-equivalence in Proposition 2.5. Since both the kernels reduce to the same group scheme over $k,$ we see that $\mathscr{M}_{1}=\mathscr{M}_{2}$ as they correspond to the Dieudonné module of the special fibre $T$. Furthermore, the tuples $(\mathscr{L}_{1},\mathscr{M}_{1})$ and $(\mathscr{L}_{2},\mathscr{M}_{2})$ are co-kernels of $(\mathscr{L},\mathscr{M})$ in the category of finite Honda systems. Therefore, Proposition 2.4 proves that $\mathscr{L}_{1}=\mathscr{L}_{2}.$ This shows that the group schemes $G_{\varphi}$ and $G_{\psi}$ are the same as subgroup-schemes of $\widetilde{\mathscr{A}}[p^{r}]$ and hence, the required kernels are the same. ∎
We end this section by proving the following proposition.
Proposition 2.7.
Let $\widetilde{A}$ and $\widetilde{B}$ be as in Lemma 2.1. Let $\widetilde{\mathscr{A}}$ and $\widetilde{\mathscr{B}}$ denote their Néron models respectively over $W(k)$ such that their special fibre is is $A.$ Then $\widetilde{A}$ and $\widetilde{B}$ are isomorphic.
Proof.
We recall the notation of $\widetilde{\alpha}$ and $\widetilde{\beta}$ mentioned before Lemma 2.6. We know that $\ker\widetilde{\alpha}=\ker\widetilde{\beta}$ by Lemma 2.6 . Therefore, we get maps $\widetilde{\varphi}:\widetilde{A}\rightarrow\widetilde{B}$ and $\widetilde{\psi}:\widetilde{A}\rightarrow\widetilde{B}$ such that $\widetilde{\varphi}\circ\widetilde{\alpha}=\widetilde{\beta}$ and $\widetilde{\psi}\circ\widetilde{\beta}=\widetilde{\alpha}$. Then we can check that $\widetilde{\varphi}\circ\widetilde{\psi}$ and $\widetilde{\psi}\circ\widetilde{\varphi}$ act as identity when we compose them with $\widetilde{\beta}$ and $\widetilde{\alpha}$ respectively, thus completing the proof of proposition. ∎
3. Local conditions for the endomorphism ring
We recall that the Newton polygon of $A[p^{\infty}]$ has slope $1/g$ and $(g-1)/g$. Moreover, we assume that the prime $p$ splits in ${\mathcal{O}}_{K}$ such that $K\otimes{\mathbb{Q}}_{p}=K^{\prime}\oplus K^{\prime}$ where $K^{\prime}$ is the unramified extensions of ${\mathbb{Q}}_{p}$ of degree $g$. We next find the local conditions of the orders that can occur as an endomorphism ring in the isogeny class of $A$ over ${\mathbb{F}}_{q}.$
3.1. Local conditions at the prime $p$
Proposition 3.1.
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomorphism algebra $K={\mathbb{Q}}(\alpha)$ where $\alpha$ is the Weil-$q$ number of $A$. Let $\operatorname{End}(A)=R.$ Then the order $R$ is maximal at $p$. More precisely, $R\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ where ${\mathcal{O}}_{K^{\prime}}$ is the maximal order in $K^{\prime}$.
Proof.
Let ${\mathbb{D}}$ denote the Dieudonné module of $A[p^{\infty}]$. Note that ${\mathbb{D}}$ admits a decomposition locally as a direct sum ${\mathbb{D}}\otimes{\mathbb{Q}}_{p}={\mathbb{D}}_{1}\times{\mathbb{D}}_{2}$ where ${\mathbb{D}}_{1}$ and ${\mathbb{D}}_{2}$ are modules over $K^{\prime}$. This further gives us a decomposition of the $p$-divisible group $A[p^{\infty}]\simeq{\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ over ${\mathbb{F}}_{q}$ where ${\mathscr{G}}_{1}$ and ${\mathscr{G}}_{2}$ have slopes $1/g$ and $(g-1)/g$ respectively (see Theorem 9.2 in [Oor08]), and ${\mathscr{G}}_{2}$ is the dual of ${\mathscr{G}}_{1}$. Since ${\mathscr{G}}_{1}$ cannot be isogenous to ${\mathscr{G}}_{2}$, we see that $\operatorname{End}({\mathscr{G}}_{1}\times{\mathscr{G}}_{2})\simeq\operatorname{End}({\mathscr{G}}_{1})\times\operatorname{End}({\mathscr{G}}_{2})$. We know that Tate’s theorem states that $R\otimes{\mathbb{Z}}_{p}\simeq\operatorname{End}A[p^{\infty}]$. Therefore, it is sufficient to prove that $\operatorname{End}({\mathscr{G}}_{1})$ and $\operatorname{End}({\mathscr{G}}_{2})$ are maximal. This argued as in [OS20], Proposition $2.1$ which we describe below. Let ${\mathscr{G}}^{\prime}$ be a $p$-divisible group isogenous to ${\mathscr{G}}_{1}$ such that $\operatorname{End}({\mathscr{G}}^{\prime})$ is maximal in its field of fractions. Let $j:{\mathscr{G}}_{0}\rightarrow{\mathscr{G}}_{1}$ be any arbitrary isogeny. Since ${\mathscr{G}}_{1}$ is isomorphic to a quotient of ${\mathscr{G}}_{0}$ by the kernel of $j$, we will show that $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves $\ker j,$ proving that $\operatorname{End}({\mathscr{G}}^{\prime})\subseteq\operatorname{End}({\mathscr{G}}_{1}),$ proving that $\operatorname{End}({\mathscr{G}}_{1})$ is maximal. Since ${\mathscr{G}}_{0}$ is a one dimensional $p$-divisible group, it has a unique subgroup of order $p^{r}$ for each $r\geq 1$. Therefore, $\operatorname{End}({\mathscr{G}}^{\prime})$ preserves the kernel of $j$. The same argument with the dual of ${\mathscr{G}}_{1}$ proves that $\operatorname{End}({\mathscr{G}}_{2})$ is maximal. ∎
3.2. Local conditions away from $p$
The above proposition proves that if $R$ is an order that occurs as an endomorphism ring of an abelian variety isogenous to $A$, then $R$ must be maximal at $p.$ Hence, it differs from the maximal order or more generally, any order that is maximal at $p$ at only finitely many primes away from $p.$ Let $l\neq p$ be any arbitrary prime. We know that since $R$ contains the Frobenius endomorphism, hence the local order ${\mathbb{Z}}_{l}[\alpha]\subseteq R\otimes{\mathbb{Z}}_{l}$ where $\alpha$ is the Weil-$q$ number of $A.$ In fact we have the following Lemma that proves that we can always adjust the local conditions of an order at $l\neq p$ up to isogeny. Thus, it guarantees the existence of an abelian variety in the isogeny class over ${\mathbb{F}}_{q}$ with endomorphism ring $R$ with the local conditions specified at primes other than $p$.
Lemma 3.2.
Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$ with endomoprhism algebra $K={\mathbb{Q}}(\alpha)$ as above. Let $R$ be an order in $K$ containing ${\mathbb{Z}}[\alpha].$ Then we can find an abelian variety $A^{\prime}$ isogenous to $A$ over ${\mathbb{F}}_{q}$ such that $\operatorname{End}(A^{\prime})\otimes{\mathbb{Z}}_{l}=R\otimes{\mathbb{Z}}_{l}$ for all primes $l\neq p.$
The proof is already given in [Wat69], Porism 4.3. For the sake completeness we present it here as well. We denote by $T_{l}(A)$ the Tate module of $A$ and let $V_{l}(A)=T_{l}(A)\otimes{\mathbb{Q}}_{l}.$
Proof.
Let $\operatorname{End}(A)=S$. Without loss of any generality we assume $S\subseteq R$. By Proposition 3.1, we know that $R\otimes{\mathbb{Z}}_{p}=S\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$, where we recall the notation of $K^{\prime}$ given in the beginning of the section. Hence the index of $S$ in $R$ is not divisible by $p$. Thus it is enough to consider primes $l\neq p$. For primes larger than the index of $S$ in $R$, the two orders are locally the same. Let $l_{1},\dots,l_{m}$ be finitely many primes where the two orders are not the same. According to Tate’s theorem $S\otimes{\mathbb{Z}}_{l}$ takes the lattice $T_{l}(A)$ to $T_{l}(A)$ in $V_{l}(A)$ for all primes $l.$ Since $K$ is commutative, $V_{l}(A)$ is a rank one module over $K\otimes{\mathbb{Q}}_{l}.$ Therefore, for any order in $K$, there exists a lattice with that prescribed order as its endomorphism ring. Let $T_{l_{1}}$ be a sub-lattice of $T_{l}(A)$ that is preserved by the order $R\otimes{\mathbb{Z}}_{l_{1}}$. The quotient of the two lattices corresponds to a finite index subgroup of $A.$ We quotient out $A$ by it to get an abelian variety $A^{\prime}$ isogenous to $A$ such that $T_{l_{1}}(A^{\prime})=T_{l_{1}}(A).$ Now repeat this process for all the other primes to finish the proof. ∎
Now for each positive integer $n$, in order get a lower bound on $I(A,{\mathbb{F}}_{q^{n}})$, we count the isomorphism classes of abelian varieties in $I(A,{\mathbb{F}}_{q^{n}})$ with endomorphism ring $R_{n}$ that is characterized by the local conditions summarised below. We note that the order $R_{n}$ is the “smallest possible” at the primes $l\neq p$ and satisfies the necessary condition at the prime $p$ as proved in Proposition 3.1.
Definition 3.3.
Let $n$ be a positive integer. We define $R_{n}\subset K={\mathbb{Q}}(\alpha)$ to be the order satisfying the following conditions:
  1. (1)
    The order $R_{n}$ contains the order ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$
  2. (2)
    It is maximal at $p,$ that is, $R_{n}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$
  3. (3)
    At the prime $l,$ we have $R_{n}\otimes{\mathbb{Z}}_{l}={\mathbb{Z}}_{l}[\alpha^{n}].$
4. Properties of abelian varieties of Lubin-Tate type.
In this section, we note certain useful properties of abelian varieties with endomorphism ring characterised in Definition 3.3 that will be useful for us in subsequent sections.
4.1. Field of definition
Lemma 4.1.
(First property) Let $B$ be an abelian variety over $k$. Let $n$ be a positive integer and let $R_{n}$ be its the endomorphism ring as defined in Definition 3.3. Then the abelian variety $B$ is defined over the field extension ${\mathbb{F}}_{q^{n}}$ of ${\mathbb{F}}_{q}$.
Proof.
Without any loss of generality, we prove this result for $n=1.$ We write $R_{1}=R.$ We can assume that $B$ is isogenous to an abelian variety $B^{\prime}$ that has endomoprhism by the maximal order so that $R\subset\operatorname{End}B^{\prime}$. Let $\varphi:B^{\prime}\rightarrow B$ be such an isogeny. Therefore the kernel of $\varphi$ is invariant under $R$, in particular it is invariant under the action of $\alpha_{B}$, the $q$-Frobenius of $B$. This implies that kernel is also Galois invariant under $\operatorname{Gal}(k/{\mathbb{F}}_{q})$ because it acts by the Frobenius of $B$. This implies that $B$ is in fact defined over ${\mathbb{F}}_{q}.$
4.2. Central leaf
Let $D_{k}$ denote the ring $W(k)[F,V]$ and $\mathbb{D}(-)$ denote the contravariant Dieudonné functor. For any positive integers $a$ and $b$, we write ${\mathscr{G}}_{a,b}$ to be the $p$-divisible group over $k$ corresponding to the Dieudonné module $\mathbb{D}({\mathscr{G}}_{a,b}):=D_{k}/D_{k}(F^{a}-V^{b})$, under the Dieudonné theory. We note that under this characterisation, the $p$-divisible group ${\mathscr{G}}_{a,b}$ will have dimension $a$ and height $a+b.$ The Dieudonné-Manin classification tells us that every $p$-divisible group over $k$ is isogenous to a product of ${\mathscr{G}}_{a,b}$ as $a$ and $b$ vary over positive integers. In fact, in dimension one we have the following stronger uniqueness result.
Lemma 4.2.
For any positive integer $h$, there exists a unique $p$-divisible group of dimension one and height $h$ over $k$. Consequently, by considering its dual the same is true for a $p$-divisible group of co-dimension one and height $h$.
Proof.
This result is well known. Let ${\mathscr{G}}$ be a $p$-divisible group of dimension one and height $h$ over $k$. let ${\mathbb{D}}({\mathscr{G}})$ be its Dieudonné module. Using the Dieudonné-Manin classification, we know that ${\mathscr{G}}$ is isogenous to ${\mathscr{G}}_{1,h-1}$ over $k$. We need to show that in fact, they are isomorphic. We first note that ${\mathbb{D}}({\mathscr{G}})$ and ${\mathbb{D}}({\mathscr{G}}_{1,h-1})$ have rank $h$ as $W(k)$-modules. We only need to show that they are isomorphic as $D_{k}$-modules. We show this by constructing a suitable $D_{k}$ bases for ${\mathbb{D}}({\mathscr{G}})$. Note that ${\mathbb{D}}({\mathscr{G}})/F({\mathbb{D}}({\mathscr{G}}))$ is a one dimensional vector space over $k$ and therefore, spanned by (the image of) an element, say $e_{1}$ in ${\mathbb{D}}({\mathscr{G}}).$ Now, let $Fe_{1}=e_{2}.$ We note that $Ve_{2}=pe_{1}.$ Next, we let $Fe_{2}=e_{3}$. Note that $e_{3}$ cannot be in the span of $e_{2}$, otherwise $e_{2}$ would be in the span of $e_{1}.$ Moreover, $e_{3}$ cannot lie in the $W(k)$-span of $e_{1}$ and $e_{2}.$ For if $e_{3}=Fe_{2}=ae_{1}+be_{2}=ae_{1}+bFe_{1}$, then $ae_{1}=0$ in ${\mathbb{D}}({\mathscr{G}})/F({\mathbb{D}}({\mathscr{G}}))$. This implies that $a=0$ which is not possible. We carry on the same process until we hit a bases such that $F$ acts as $Fe_{i}=e_{i+1}$ for $11\leq i\leq h-$ 1. Finally writing $Fe_{h}=a_{1}e_{1}+a_{2}e_{2}+\dots+a_{h}e_{h}$ and applying $V$ on both sides implies that $pe_{h}=a_{1}Ve_{1}$ and $a_{i}=0$ for $i\geq 2$. Since $e_{1}$ is not in the image of $F,$ we see that $a_{1}=p$ and hence $Fe_{h}=pe_{1}$. This gives us an isomorphism between ${\mathscr{G}}$ and ${\mathscr{G}}_{1,h-1}$.
Definition 4.3.
(Central leaves) Let ${\mathscr{G}}$ be a $p$-divisible group over ${\mathbb{F}}_{q}.$ Then the ${\mathbb{F}}_{q^{n}}$ points of the central leaf of ${\mathscr{G}}$ in $\mathcal{A}_{g,1}\otimes{\mathbb{F}}_{q}$ is defined as
$c({\mathscr{G}})({\mathbb{F}}_{q^{n}})=\{B\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})\mid{\mathscr{G}}\otimes k\simeq B[p^{\infty}]\otimes k\}.$
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a principally polarised abelian variety. Then the central leaf $c(A)$ is defined to be the central leaf associated to its $p$-divisible group $A[p^{\infty}]$.
We note that $c(A)$ in fact cuts out a closed sub-variety in the Newton stratum of $A$.
Lemma 4.4.
(Second property) Let $n$ be a positive integer. Let $B$ be an abelian variety over ${\mathbb{F}}_{q^{n}}$ satisfying the conditions mentioned at the beginning of Section 3. Let $\operatorname{End}(B)=R_{n}\subset K$. Then $B$ lies in the central leaf of the $p$-divisible group ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$.
Proof.
By applying Proposition 3.1 to the abelian variety $B$, we see that $\operatorname{End}(B[p^{\infty}])\otimes k={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. This implies that its Dieudonné module ${\mathbb{D}}\otimes k$ splits into a direct sum ${\mathbb{D}}_{1}\oplus{\mathbb{D}}_{2}$ such that $\operatorname{End}{\mathbb{D}}_{1}$ and $\operatorname{End}{\mathbb{D}}_{2}$ are maximal at $p.$ This gives the desired decomposition of $B[p^{\infty}]\otimes k$ into a product ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$. By Dieudonné-Manin classification ${\mathscr{G}}_{1}\times{\mathscr{G}}_{2}$ is isogenous to ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$. Let $\psi$ be such an isogeny. Since the endomorphism ring preserves the kernel of $\psi$, it must also factor into a direct summand. This shows that ${\mathscr{G}}_{1}$ (respectively ${\mathscr{G}}_{2}$) is isogenous hence to ${\mathscr{G}}_{1,g-1}$ (respectively ${\mathscr{G}}_{g-1,1}$). By Lemma 4.4, we see that they’re in fact isomorphic. This shows that the isogeny $\psi$ is an isomorphism and hence $B$ lies in the central leaf of ${\mathscr{G}}_{1,g-1}\times{\mathscr{G}}_{g-1,1}$. ∎
5. Isogeny classes and the class group
Let $R\subset K$ be any order. Let $I\subset K$ be an invertible $R$ module. That is, $I$ is a fractional ideal of $R_{n}$ such that there exists some fractional ideal $J$ with $IJ=R$. We say two fractional ideals, $I_{1}$ and $I_{2}$ are equivalent if and only if there exists some $a\in K^{\times}$ such that $I_{1}=aI_{2}.$
Definition 5.1.
(Class group of an order) Under the equivalence mentioned above, we define the equivalence classes of fractional ideals of $R$ to be the class group of $R$, denoted as $Cl(R).$
The heart of rest of the paper lies in the following proposition that gives us a convenient characterisation of (isogenous) abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$ in Definition 3.3. Here we recall the setup and notation at the beginning of Section 2. In particular, we return to the set up in characteristic zero with $\widetilde{A}$ an abelian variety over $L=W({\mathbb{F}}_{q})(1/p)$. Let $\Phi$ be the primitive CM type on $K$ corresponding $\widetilde{A}$ over $L$. We denote by $A$ over ${\mathbb{F}}_{q}$ its reduction mod $p$.
Proposition 5.2.
For any positive integer $n$, there exists an injective map from the class group of $R_{n}$ to the set of isomorphism classes of abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}\subset(K,\Phi).$
Proof.
Let $I$ be an element of $Cl(R_{n}).$ We recall that $\Phi$ gives is an isomorphism $K\otimes{\mathbb{C}}={\mathbb{C}}^{\Phi}$ where ${\mathbb{C}}^{\Phi}$ denotes $g$ copies of ${\mathbb{C}}$ indexed by the embeddings in $\Phi.$ Since $I$ is a lattice in $K,$ we can embed it inside ${\mathbb{C}}^{\Phi}$. We refer to this embedding as $\Phi(I).$ The complex torus ${\mathbb{C}}^{\Phi}/\Phi(I)$ admits a polarisation (see Section 7) and therefore we get an abelian variety over ${\mathbb{C}}$ which we denote by $A_{\Phi(I)}$. Since $\Phi$ is primitive, $A_{\Phi(I)}$ is simple and so its endomorphism algebra is $K$. By construction $\operatorname{End}(A_{\Phi(I)})=R_{n}.$ Furthermore, since $A_{\Phi(I)}$ is a CM abelian variety, it descends, with its entire endomorphism algebra to a number field. Now, if the abelian variety had CM by the maximal order, then the abelian variety is defined over the maximal unramified extension of the reflex field $K^{*}$ and hence it would be defined over an unramified extension of ${\mathbb{Q}}_{p}.$ In our case, however, the field of definition might be slightly larger than that and may be a ramified extension of ${\mathbb{Q}}_{p}$. We first rule out this possibility for the abelian variety $A_{\Phi(R_{n})}.$ We note that $A_{\Phi(R_{n})}$ is isogenous to the abelian variety $A^{\prime}$ corresponding to the lattice ${\mathcal{O}}_{K}$ in $K$. Now, $A^{\prime}$ has CM by a maximal order over ¯$\overline{{\mathbb{Q}}}_{p}.$ Let $F$ be the unramified extension of ${\mathbb{Q}}_{p}$ where $A^{\prime}$ is defined. Since the kernel of the isogeny has order equal to the index of $R_{n}$ in ${\mathcal{O}}_{K}$, we can embed it inside $A^{\prime}[m]$ for some $(m,p)=1$. Now the kernel is defined over some extension $F^{\prime}$ of $F$. This must be unramified because the inertia group $I(\overline{v})$ acts trivially on the $m$-torsion and therefore on the subgroup as well. Here $\bar{v}$ refers to the extension of a place $v$ of $F^{\prime}$ to its separable closure. This descends $A_{\Phi(R_{n})}$ to an unramified extension of ${\mathbb{Q}}_{p}.$ This same argument works in descending an abelian variety corresponding to a fractional ideal $I$ whose index in $R_{n}$ is prime to $p.$ So suppose that $P$ is a fractional ideal of $R_{n}$ with index a power of $p$. We extend this to a fractional ideal $P^{\prime}$ of ${\mathcal{O}}_{K}$. Since $R_{n}$ is maximal at $p$, the index of $P$ in $R_{n}$ equals the index of $P^{\prime}$ in ${\mathcal{O}}_{K}$. Hence the index of $P$ in $P^{\prime}$ is prime to $p.$ Since the abelian variety corresponding to $P^{\prime}$ has endomorphism by the maximal order, it descends to an unramified extension of ${\mathbb{Q}}_{p}.$ We can now repeat above argument with $A_{\Phi(P)}$. We note that all the abelian varieties constructed above are isogenous to $\widetilde{A}_{L^{\prime}}$ over $L^{\prime}$ by Lemma 2.1. Finally, under the equivalence of ideals in $Cl(R_{n})$, two equivalent ideals would correspond to isomorphic lattices and hence give us isomorphic abelian varieties. ∎
Corollary 5.3.
Let $n$ be any positive integer. Then there exists an injective map from the class group $Cl(R_{n})$ to the set of isomorphism classes of abelian varieties over ${\mathbb{F}}_{q^{n}}$ isogenous to $A$ (without any condition on polarisation).
Proof.
By Proposition 2.7, the reduction mod $p$ map from the set of isomorphism classes of abelian varieties isogenous to $\widetilde{A}$ over $L^{\prime}$ to that of $A$ over $k$ is injective. Moreover, for any $n$, by Lemma 4.1 the abelian variety constructed using an ideal $I\subset R_{n}$ is defined over $W({\mathbb{F}}_{q^{n}})(1/p)$ and hence the reduction mod $p$ of their Néron models descend to ${\mathbb{F}}_{q^{n}}$. By Lemma 2.2, their endomorphism ring is $R_{n}$ and hence have $\alpha^{n}\in K$ as their Frobenius morphism. ∎
6. Estimates
The goal of this section is to find the asymptotics for $Cl(R_{n})$. We prove the following estimate.
Proposition 6.1.
For a positive density set of positive integers $n$, the class group $Cl(R_{n})=q^{n{((g^{2}-2g)/2+o(1))}}.$
Proof.
As $n$ grows large, the class group of an order is well approximated by the square root of its discriminant. Let $\alpha_{i}$, where $1\leq i\leq 2g$ and $\alpha_{1}=\alpha$ be the conjugates of $\alpha.$ For each $i$, we write $\alpha_{i}=q^{1/2}(\cos{\theta}+i\sin{\theta}).$ Then discriminant of the monogenic order ${\mathbb{Z}}[\alpha^{n}]$ is given by
$\operatorname{disc}({\mathbb{Z}}[\alpha^{n}])=\prod_{1\leq i<j\leq 2g}({\alpha^{n}_{i}}-{\alpha^{n}_{j}})^{2}=q^{\frac{2ng(2g-1)}{2}}\prod_{1\leq i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{i})-\cos(n\theta_{j})-i\sin(n\theta_{j}))^{2}$
Since $\operatorname{disc}({\mathbb{Z}}[\alpha^{n}])=[{\mathbb{Z}}[\alpha^{n}]:R_{n}]^{2}\operatorname{disc}(R_{n})$, it is enough to estimate the index of ${\mathbb{Z}}[\alpha^{n}]$ in $R_{n}.$ We note that the index is $1$ at all primes $l\neq p$ (see Definition 3.3, point (3)), hence it is sufficient to estimate the index after tensoring the orders with ${\mathbb{Z}}_{p}$. This is done below.
6.1. Estimating the index at the prime $p$
We write ${\mathbb{Z}}[\alpha]\otimes{\mathbb{Z}}_{p}\subset{\mathcal{O}}_{K}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. Under this embedding, the image of $\alpha$ is $(\beta,\gamma)$ where $v_{{\mathfrak{q}}}(\beta)=1/g$ and $v_{q}(\gamma)=(g-1)/g$. The ring ${\mathbb{Z}}_{p}[(\beta,\gamma)]$ is monogenic and is isomorphic to ${\mathbb{Z}}_{p}[x]/(f_{1}f_{2})$ where $f=f_{1}f_{2}$ is the factorization of the minimal polynomial $f$ of $\alpha$ in ${\mathbb{Z}}_{p}.$ Let $\beta_{i}$ and $\gamma_{i}$ be the roots of $f_{1}$ and $f_{2}$ with $\beta_{1}=\beta$ and $\gamma_{1}=\gamma$ and $1\leq i\leq g$. Now, for any positive integer $n$, we are reduced to finding the index of ${\mathbb{Z}}_{p}[\alpha^{n}]={\mathbb{Z}}_{p}[(\beta^{n},\gamma^{n})]$ in ${\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$. We compute this as follows:
Consider the monogenic ring ${\mathbb{Z}}_{p}[x]/g_{1}g_{2}:={\mathbb{Z}}_{p}[\delta]$ where $\delta=(\beta/q^{1/g},\gamma/q^{g-1/g})$ and the roots of $g_{1}$ and $g_{2}$ are given by $\beta^{\prime}_{i}:=\beta_{i}/q^{1/g}$ and $\gamma^{\prime}_{i}:=\gamma_{i}/q^{g-1/g}$ respectively where $1\leq i\leq g$. Thus, we have the following containment of orders:
${\mathbb{Z}}_{p}[\alpha^{n}]\subseteq{\mathbb{Z}}_{p}[\delta^{n}]\subseteq{\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$
Now for most $n$, the index ${\mathbb{Z}}_{p}[\delta^{n}]\subseteq{\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ remains constant, that is does not depend on $n$. Indeed, the discriminant of ${\mathcal{O}}_{K^{\prime}}\oplus{\mathcal{O}}_{K^{\prime}}$ is constant, while that of ${\mathbb{Z}}_{p}[\delta^{n}]$ is
$\operatorname{disc}({\mathbb{Z}}_{p}[\delta^{n}])=\prod_{1\leq i<j\leq g}((\beta^{\prime}_{i})^{n}-(\beta^{\prime}_{j})^{n})((\gamma^{\prime}_{i})^{n}-(\gamma^{\prime}_{j})^{n})\prod_{1\leq i,j\leq g}((\beta^{\prime}_{i})^{n}-(\gamma^{\prime}_{j}))$
$={\mathbb{Z}}_{p}[\delta]\prod_{1\leq i<j\leq g}((\beta^{\prime}_{i})^{n-1}+(\beta^{\prime}_{i})^{n-2}\beta^{\prime}_{j}+\dots+(\beta^{\prime}_{1})^{n-1})((\gamma^{\prime}_{i})^{n-1}+(\gamma^{\prime}_{i})^{n-2}\gamma^{\prime}_{j}+\dots+(\gamma^{\prime}_{1})^{n-1})$
$((\gamma^{\prime}_{i})^{n-1}+(\gamma^{\prime}_{i})^{n-2}\beta^{\prime}_{j}+\dots+(\beta^{\prime}_{j})^{n-1})$
We recall that $K^{\prime}$ equals the unramified extension of ${\mathbb{Q}}_{p}$ of degree $g$ and hence, for $n$ divisible by $p^{g}-1$ then all $\beta_{i}$ and $\gamma_{j}$ with $1\leq i,j\leq g$ are congruent to $1\operatorname{mod}$ $p$. Moreover, if we take $n$ co-prime to $p$, then the discriminant of ${\mathbb{Z}}_{p}[\beta^{n}]$ and ${\mathbb{Z}}_{p}[\beta]$ differ by a unit.
Hence, for such $n$ it is enough to compute the index of ${\mathbb{Z}}[\alpha^{n}]$ in ${\mathbb{Z}}[\delta^{n}]$. This is is straightforward using the relation:
$\operatorname{disc}{\mathbb{Z}}_{p}[\delta^{n}]\cdot[{\mathbb{Z}}_{p}[\alpha^{n}]:{\mathbb{Z}}_{p}[\delta^{n}]]^{2}=\operatorname{disc}{\mathbb{Z}}_{p}[\alpha^{n}]$
The index is given by the following expression for $n=1$ and can be raised to the power $n$ for any positive integer $n$.
$[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]=\frac{\prod_{1\leq i<j\leq g}(\beta_{i}-\beta_{j})(\gamma_{i}-\gamma_{j})\prod_{1\leq i,j\leq g}(\beta_{i}-\gamma_{j})}{\prod_{1\leq i<j\leq g}(\beta_{i}/q^{1/g}-\beta_{j}/q^{1/g})(\gamma_{i}/q^{(g-1)/g}-\gamma_{j}/q^{(g-1)/g)})\prod_{1\leq i,j\leq g}(\beta_{i}/q^{1/g}-\gamma_{j}/q^{(g-1)/g})}$
which equals $(q^{1/g})^{g(g-1)/2}\cdot(q^{(g-1)/g})^{g(g-1)/2}\cdot(q^{1/g})^{g^{2}}=q^{(g^{2}+g)/2}$ up to some constant. Hence, for a positive density set of positive integers, we get $[{\mathbb{Z}}[\alpha^{n}]:R_{n}]=[{\mathbb{Z}}_{p}[\alpha]:{\mathbb{Z}}_{p}[\beta]]=q^{n(g^{2}+g)/2}$ up to some constant.
6.2. Final estimate
For a positive density set of positive integers $n$, the discriminant of $R_{n}$ equals (up to some constant)
$\frac{(\operatorname{disc}({\mathbb{Z}}[\alpha^{n}]))^{1/2}}{[{\mathbb{Z}}[\alpha^{n}]:R_{n}]}=\frac{q^{\frac{ng(2g-1)}{2}}\prod_{i<j\leq 2g}(\cos(n\theta_{i})+i\sin(n\theta_{j})-\cos(n\theta_{j})-\sin(n\theta_{j}))^{2}}{q^{n(g^{2}+g)/2}}$
Since the term with sines and cosines is bounded, we see that the class group $Cl(R_{n})$ is well approximated by the leading term $q^{\frac{ng(2g-1)}{2}-\frac{n(g^{2}+g)}{2}}=q^{n(g^{2}-2g)/2}$. ∎
Corollary 6.2.
For a positive density set of positive integers $n$, the size of isogeny class of $A$ (without any conditions on polarisation) in characteristic zero is greater than or equal to $q^{n(\frac{(g^{2}-2g)}{2}+o(1))}$.
Proof.
This follows from Corollary 5.3 and the calculation above in Proposition 6.1. ∎
7. Polarisation
So far, we have an estimate for the size of the isogeny class without specifying any condition on the polarisation. In order to restrict this count to the isogeny class in $\mathcal{A}_{g,1},$ we need to impose further conditions for principal polarisation. In [How95], Howe describes the polarisation data for ordinary abelian varieties which is summed up nicely in [ST18], Section 3.1. However, since the data is described purely in characteristic zero and is independent of the type of abelian variety, we use the same description in our case that we describe below.
7.1. Polarisation data.
We work with the CM type $\Phi$ on $K$ as before. For any element $x\in K$, we denote $\overline{x}$ to be the complex conjugation (induced by $K$) on $x$. The polarisation on an abelian variety corresponding to the ideal $I$ in $Cl(R_{n})$ (see Proposition 5.2) is given by a ${\mathbb{Q}}$-valued bilinear form on $K$.
$\Psi:(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(ux\overline{y})$
satisfying the following properties.
  1. (1)
    The form $\Psi$ restricted to $I$ is valued in ${\mathbb{Z}}$.
  2. (2)
    The element $u\in K$ is purely imaginary.
  3. (3)
    For all $\varphi\in\Phi$, we have $\varphi(u)/i$ is positive.
  4. (4)
    The dual of an ideal $I$ is the trace dual of $\overline{I}$ where $\overline{I}$ is the CM conjugate of $I$ with respect to the unique involution on $K.$
We note that if an ideal $I$ is dual to itself with respect to the above form then then its corresponding abelian variety will be principally polarized. We have the following proposition.
Proposition 7.1.
Let $R_{n}^{+}$ denote the intersection of $R_{n}$ with the maximal totally real field $K^{+}\subset K.$ Then the kernel of the norm map
$N:Cl(R_{n})\rightarrow Cl^{+}(R_{n}^{+})$
is either empty or is in one-to-one correspondence with principally polarised abelian varieties isogenous to $\widetilde{A}$ over $L^{\prime}$ with endomorphism ring $R_{n}$ (and hence isogenous to $A$ over ${\mathbb{F}}_{q^{n}}$). Here $Cl^{+}(R_{n}^{+})$ denotes the narrow class group of the order $R_{n}^{+}$ and the norm map refers to $I\mapsto I\overline{I}$.
Proof.
The proof of this proposition goes through verbatum as in Lemma 3.8 in [ST18]. ∎
Lemma 7.2.
For a positive density set of integers $n,$ we have $\frac{|Cl(R_{n})|}{|Cl^{+}(R_{n}^{+})|}=q^{(n(g+1)(g-2)/2+o(1))}$
Proof.
It suffices to find an upper bound for $Cl^{+}(R_{n}^{+})$. Let $\lambda^{(n)}$ be the element $\alpha^{n}+q^{n}/\alpha^{n}$. We denote the $p$-adic conjugates of $\lambda^{(n)}$ by $\lambda^{(n)}_{i}$ for $1\leq i\leq g$ and $\lambda^{(n)}_{1}:=\lambda^{(1)}$. As before, in order to estimate the class group, we first estimate the index $[{\mathbb{Z}}[\lambda^{(n)}]:R_{n}^{+}].$ For simplicity we do the computation with $n=1$ as for a general $n$, the estimate is raised to the power of $n$. We write $\lambda^{(1)}_{i}=\lambda_{i}$ for $1\leq i.j\leq g$ and $R^{+}_{1}:=R^{+}$. It is enough to compute the index of ${\mathbb{Z}}[\lambda]$ in $R^{+}$ at the prime $p$. We recall $K^{\prime}$ is the unramfied extension of ${\mathbb{Q}}_{p}$ of degree $g$. As $R^{+}\otimes{\mathbb{Z}}_{p}={\mathcal{O}}_{K^{\prime}}$, we can follow the the same argument as in the proof of Proposition 6.1 to compute this. Hence, for a positive proportion of positive integers $n$ (satisfying the same congruent conditions as in the proof of Proposition 6.1), the index is given by the following expression:
$[{\mathbb{Z}}[\lambda]:R^{+}]=\frac{\prod_{1\leq i<j\leq g}\lambda_{i}-\lambda_{j}}{\prod_{1\leq i<j\leq g}(\lambda_{i}/q^{1/g}-\lambda_{j}/q^{1/g})}=q^{\frac{{g-1}}{2}}$
The discriminant of the monogenic ring ${\mathbb{Z}}[\lambda]$ is calculated easily, and is also given in Lemma 3.8 in [ST18] where the leading the term is given by $q^{g(g-1)/2}.$ Now, the leading term of the $\operatorname{disc}(R^{+})$ is approximated by dividing this discriminant by the square of the index computed above. Hence $Cl(R^{+})$ comes out to be $q^{(g^{2}-3g+2)/4}$, that is taking the square root of $\operatorname{disc}(R^{+})$. Combining this (raised to the power of $n$) with the estimate of $Cl(R_{n})$ in Proposition 6.1 we get the desired result. ∎
We note that a polarisation on $\widetilde{A}$ extends to a polarisation on the Néron model and hence, reduces to a polarisation on its special fibre $A$ as well. Therefore, the count in characteristic zero restricted to principally polarised varieties gives us the right estimate in characteristic $p$. It remains to prove that the kernel of the Norm map in Proposition 7.1 is non empty. That is, the subset of $I(\widetilde{A},L^{\prime})$ with endomorphism ring $R_{n}$ is not empty. This is proved in the final proposition of this paper.
Proposition 7.3.
We assume that $q^{1/2},\alpha=\alpha_{1},$ and its conjugates $\alpha_{2},\dots\alpha_{g}$, are multiplicatively independent. Then for a positive density set of integers $n$, there exist principally polarised abelian varieties over $L^{\prime}$ with endomorphism ring $R_{n}$.
Proof.
Let $h_{n}$ denote the minimal polynomial of $\alpha^{n}+(q/\alpha)^{n}$ and $h_{n}^{\prime}$ be its derivative. Let $u_{n}=(\alpha^{n}-(q/\alpha)^{n}h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n}))^{-1}$. We note that $\alpha^{n}-(q/\alpha)^{n}$ is purely imaginary while $h^{\prime}_{n}(\alpha^{n}+(q/\alpha)^{n})$ is purely real. Hence $u_{n}$ is purely imaginary for all $n$. By (the proof of) Proposition 9.4 in [How95], the trace dual of ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ is $u_{n}^{-1}{\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$. Therefore, if we consider the bilinear form
$(x,y)\mapsto\operatorname{tr}_{K/{\mathbb{Q}}}(u_{n}x\overline{y})$
then the dual of ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ with respect to this form is itself. Furthermore, the above form induces a polarisation on the elements of $Cl(R_{n})$ when $\varphi(u_{n})/i$ is a positive real number for every $\varphi\in\Phi.$ The argument given in Proposition $3.6$ of [ST18] shows that for a positive proportion of $n$, $\varphi(u_{n})/i$ is positive under the assumption that the set $\{q^{1/2},\alpha_{1},\dots,\alpha_{g}\}$ is multiplicatively independent. We assume that $n$ is an integer such that $u_{n}$ satisfies appropriate conditions so that we get a polarisation on the elements of $Cl(R_{n}).$ Without loss of any generality, we assume $n=1$ as the argument remains the same for any $n\geq 1$. Let $\widetilde{A}/L^{\prime}$ be the abelian variety corresponding to the order $R_{1}:=R$ and let $A$ be its reduction mod $p$ over ${\mathbb{F}}_{q}$. Further, let ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2}$ be the intersection of prime ideals in ${\mathcal{O}}_{K}$ lying over $p$ with $R$. Since $R$ is locally maximal at the primes lying over $p$, the prime ideals ${\mathfrak{p}}_{1}$ and $\mathfrak{p}_{2}$ are invertible. Note that ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and $R$ agree everywhere except at ${\mathfrak{p}}_{1}$ and ${\mathfrak{p}}_{2},$ and hence the same is true for ${\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]^{\vee}={\mathbb{Z}}[\alpha^{n},(q/\alpha)^{n}]$ and $R^{\vee}$. Therefore, $R$ is ${\mathfrak{p}}_{1}^{m}{\mathfrak{p}}_{2}^{k}$, while $R^{\vee}={\mathfrak{p}}_{1}^{k}{\mathfrak{p}}_{2}^{m}$ for some integers $m$ and $k.$ Now, the lattice $R$ locally admits a decomposition corresponding to the slope $1/g$ and $g-1/g$ of $A$. We write this decomposition as $T_{1/g}\oplus T_{g-1/g}.$ We recall that $T_{1/g}$ and $T_{g-1/g}$ have endomorphism by the maximal order in $K^{\prime}$ where $K^{\prime}$ is the unramified extension of ${\mathbb{Q}}_{p}$ of degree $g$. Hence, the kernel of the corresponding polarisation map $\widetilde{A}\rightarrow\widetilde{A}^{\vee}$ corresponds to the direct sum of $\overline{\omega}^{m}$-torsion of $T_{1/g}$ and $\overline{\omega}^{k}$-torsion of $T_{g-1/g}$ where $\overline{\omega}$ is the uniformizer of ${\mathcal{O}}_{K^{\prime}}$. Since the polarisation map has a kernel that is self dual, we see that $m=k.$ Therefore, the kernel equals to the $p^{gm}$-torsion of $\widetilde{A}$ and hence $\widetilde{A}$ is principally polarised as $\widetilde{A}^{\vee}\simeq\widetilde{A}/\widetilde{A}[p^{gm}]\simeq\widetilde{A}$. ∎